### v2

# Newcastle-upon-Tyne, UK

Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational assumption that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple application contexts. Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher. We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities. These results call into question the widely held assumption that adding more sensors inherently yields higher security. We ultimately caution against relying on raw sensor data as a primary source of randomness.
Keywords:
Entropy Sensors Mobile security
1 Introduction
Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This rich sensor data has enabled applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while newer techniques incorporate ambient phenomena—characteristics of magnetic fields [2] and thermal fluctuations [3]—to mitigate relay attacks or reduce user-driven authentication prompts. Despite these advancements, fundamental questions remain. Multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 5, 6], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 6, 7, 2, 5], that do not address core cryptographic requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
Our analysis reveals systemic limitations. Commodity sensors exhibit significant biases, with gyroscopes yielding min-entropy as low as 0.96 bits—well below cryptographic thresholds. Although multi-modal sensor fusion mitigates some of these issues, non-uniform distributions and inter-sensor correlations reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy. These findings challenge the notion that increasing the number of sensors automatically strengthens security, and underscore the inadequacy of raw sensor data as a reliable entropy source. Our contributions are:
  • We introduce the first systematic approach to evaluating sensor entropy across diverse modalities and datasets using various entropy metrics (max, Shannon, collision, and min-entropy).
  • We empirically demonstrate how inter-sensor correlations and biases erode entropy, casting doubt on the proposition that using multiple sensors adds substantially to security.
  • Ultimately, we advise against relying on commodity sensors as primary sources of randomness for security-critical applications, both on a single- and multiple-sensor basis.
The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis framework is made publicly available to foster future research. [^1] https://github.com/cgshep/entropy-collapse-mobile-sensors
2 Background
This section discusses sensor-based applications primitives, critiques existing approaches, and formalises the entropy metrics underpinning our analysis.
2.1 Mobile Sensor-based Security Applications
One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [8, 6, 7, 4]. Mehrnezhad et al.[6] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together. The authors posit that “physical tapping causes random but correlated vibrations at both devices, which are hard to forge (or reproduce)” (p.1, [6]). The work reports an equal error rate (EER) of 17.65% using a machine learning-based approach. Gurulian et al. [7] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device. Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [9, 10, 11, 12, 13, 14, 15, 16], where sensor data is used to authenticate users passively without explicit interaction. (with up to 99% classification accuracy in some instances [10]). Mobile sensors have also been proposed as enabling novel device pairing schemes for secure communication, e.g. [1, 17].
The notion of ‘hardness’ is typically inferred based on a machine learning-based analysis; it is not grounded in standard quantitative methods used in the field of authentication, such as Shannon [18] and min-entropy [19]. A general approach to follows one of those given by Shrestha et al. [20]. Here, sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 6]. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [12], EERs [7, 6], Receiver-Operator Curves (ROC) [16], and accuracy [16] have been used to evaluate the effectiveness of distinguishing between legitimate and illegitimate samples. User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [11, 1]. However, machine learning-based evaluation metrics do not measure intrinsic entropy—a crucial factor in determining whether sensor data is a suitable source of randomness for security applications. Some work has attempted to assess the entropy of sensor signals, but this represents a minority. T2Pair by Li et al. [21] 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al. [22] reporting 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards.
We point to adjacent work that has questioned the utility of mobile sensors. Markantonakis et al. [4], building on Gurulian et al. [23] and Shepherd et al. [24], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints. Using consumer off-the-shelf devices, the work reports that sensors available through the Android platform could not be used to reliably detect man-in-the-middle attacks in NFC contactless transactions within the 500ms time window required by the EMV payment protocol. Under those constraints, the effectiveness drastically decreased compared to previously reported results. Indeed, it was lower to the extent such that sensors could not be recommended for mobile transactions without posing significant usability and security issues in practice.
2.2 Sensor Entropy Analyses
In earlier work, Voris et al. [19] investigated the use of accelerometers as true random number generators (TRNGs), specifically aboard a WISP RFID tag and a Nokia N97 mobile phone. The authors find that min-entropy—defined in §2.3—that min-entropy is proportional tot he motion applied to the device, and that stationary movement has the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC) are considered other significant influences on the generated entropy. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [25] analyse three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method used. Krhovják et al. [3] examine the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropy values of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [26] present an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. The work uses a single-sensor analysis, presenting min-entropy values between 0–7.85. Motion sensors, e.g. accelerometer and vibration sensors, yielded the highest entropy.
Mobile sensors have also been proposed as entropy sources in low-cost RNG designs. Suciu et al. [27] propose using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values are presented. Wallace et al. [29] explore the use of accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Mobile sensors are inherently opportunistic entropy sources, contrasting with dedicated hardware like ring oscillators, those build on Johnson-Nyquist thermal noise, or quantum-based generators (e.g. see [30]). Mobile sensors depend heavily on user behaviour and environmental, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic assessments or those based on machine learning-based evaluation metrics [12, 16]. Such approaches do not account well for skewed distributions and other biases in the underlying data. This contrasts with typical entropy measures used in RNG analyses or in the study of authentication systems [31]. For instance, NIST SP800-90B [32] and AIS 20/31 [33] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
2.3 Definitions
We use the following definitions and notation throughout this work.
Definition 1 (Rényi Entropy)
Let $X$ be a discrete random variable taking values in a set $\mathcal{X}$ with probability mass function $p(x)$. The Rényi entropy of order $\alpha$ ($\alpha>0$, $\alpha\neq 1$) is defined as
$H_{\alpha}(X)\,=\,\frac{1}{1-\alpha}\,\log\!\Biggl{(}\sum_{x\in\mathcal{X}}p(x)^{\alpha}\Biggr{)}.$ (1)
We draw attention to four special cases of $\alpha$ that are widely used in the literature. Firstly, the Hartley (Max) Entropy ($\alpha=0$), given in Eq. 2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
$H_{0}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
Second is the Shannon Entropy (Eq. 3), which corresponds to the classical definition of entropy in information theory, and is the limit of $H_{\alpha}$ as $\alpha\to 1$.
$H_{1}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
Another case is the Collision Entropy ($\alpha=2$), quantifying the probability of “collisions” of multiple draws from $X$. This is given in Eq. 4.
$H_{2}(X)\,=\,-\log\Bigl{(}\sum_{x\in\mathcal{X}}p(x)^{2}\Bigr{)}.$ (4)
$H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy is thus used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes. This is the Min-Entropy (Eq. 5), i.e. the value at the limit $\alpha\to\infty$.
$H_{\infty}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
Note that $H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq H_{2}(X)\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in NIST SP800-90B [32] and AIS 20/31 [33]. We also rely on joint entropy for assessing the entropy of multiple random variables.
Definition 2 (Joint Rényi Entropy)
Let $X_{1},X_{2},\dots,X_{n}$ be discrete random variables that jointly take values in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$, with joint probability mass function $p(x_{1},x_{2},\dots,x_{n})$. The Rényi entropy of order $\alpha$ ($\alpha>0,\alpha\neq 1$) for these $n$ variables is defined as the following, where the sum is taken over all $(x_{1},\dots,x_{n})$ in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$:
$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},\dots,x_{n})}p(x_{1},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
In the limit $\alpha\to 1$, then $H_{\alpha}(X_{1},X_{2},\dots,X_{n})$ converges to the classical joint Shannon entropy of these $n$ variables. To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we also define the Kullback-Leibler divergence.
Definition 3 (Kullback-Leibler (KL) Divergence)
Given a true probability distribution $P(x)$ of a random variable and an approximate or reference distribution $Q(x)$, the KL divergence is defined as follows:
$D_{KL}(P\;||\;Q)=\sum_{x\in\mathcal{X}}P(x)\log\Big{(}\frac{P(x)}{Q(x)}\Big{)}$ (7)
3 Experiment Design
3.1 Threat Model and Assumptions
We consider an adversary whose primary objective is to compromise sensor-based key generation or authentication schemes by exploiting the inherent biases or predictable patterns within sensor data. In particular, we assume the adversary has full or near-full knowledge of the probability distribution of sensor outputs. They may obtain this through statistical modeling of publicly available datasets, prolonged observation of target users, or eavesdropping on sensor data streams.
We exclude adversarial capabilities such as fault-injection attacks or hardware attacks, e.g. see [34]. While such attacks could further reduce the effective entropy space, we regard them as a higher level of sophistication—an important future research direction but not the focus here. Lastly, rather than brute-forcing all possible sensor outputs, we assume a more strategic approach: the adversary uses its knowledge of the sensor’s distribution to prioritise the most probable (or most easily predicted) outcomes. By iterating guesses in order of descending probability, they minimise their expected effort. This threat model represents a realistic adversary who can capitalise on statistical biases in sensor data but does not directly sabotage or physically compromise the device. Our goal is to evaluate whether typical sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist such knowledge-driven attacks.
3.2 High-level Methodology
We aim to determine the global (i.e., population-level) entropy characteristics of various mobile sensors under ordinary usage conditions, rather than focusing on per-user or scenario-specific differences. This choice reflects common real-world deployments, which often must accommodate a wide range of behaviors and environments with minimal tailoring. Our approach involves five main stages:
  1. 1.
    We acquire large-scale sensor readings from publicly available datasets that capture diverse user activities and device usage patterns. These datasets encompass different motion, environmental, and orientation sensors. Detailed descriptions of each dataset are provided at the end of this section.
  2. 2.
    We merge sensor readings into a single, global distribution for each sensor modality in each dataset. For sensors that are inherently discrete or quantised (e.g., integer output ranges), we simply count occurrences. For sensors that produce (quasi-)continuous values, we rely on quantisation strategies (Freedman–Diaconis binning) to partition the output space and approximate an empirical probability mass function.
  3. 3.
    From these global distributions, we compute max, Shannon and collusion entropies to measure the best- and average-case uncertainties of sensor outputs, along with the min-entropy to characterise the worst-case unpredictability.
  4. 4.
    Many real-world proposals combine multiple sensor streams to purportedly increase security. To assess the impact on worst-case unpredictability, we use Chow-Liu trees to approximate the joint distributions of different sensor modalities. This allows us to estimate higher-dimensional entropies without incurring prohibitive computational costs. We discuss this in §4.3.
  5. 5.
    Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary. We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to similar predictability concerns.
Throughout this process, we remain mindful of well-documented constraints with NIST SP 800-90B, SP 800-22 [32], and AIS 20/31 [33] in analysing multi-sensor data streams [35, 25]. Such frameworks were not designed to analyse the joint entropy of complex, multivariate data sources, which is the aim of this work. (For example, NIST SP 800-90B focusses on assessing univariate entropy sources with reduced, i.e. 8-bit, output sizes [35, 32]). To begin with, we sought publicly available sensor datasets suitable for analysing motion and environmental data at scale. Our search involved broad queries across IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub. Several ostensibly “open” datasets either were no longer downloadable or imposed restrictive licensing terms [36, 37, 38]. Ultimately, we narrowed our scope to four datasets that offer diverse usage contexts, consistent sampling rates, and documented sensor modalities:
  • UCI-HAR [39]: A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities. Data includes triaxial accelerometer and gyroscope signals.
  • University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz. Multiple Huawei Mate 9 phones attached across the users’ body in a body area network (BAN). The publicly available SHL Preview dataset is used, comprising three recording-days per user (a total of 59 hours of data). To scope this study, we use the motion sensors dataset from the handheld mobile phone as a good fit with related work (see §2).
  • Relay [23]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés). The dataset encompasses accelerometer, gyroscope, and environmental readings taken in realistic payment scenarios.
  • PerilZIS [42]: Collected at 10 Hz from a Texas Instruments SensorTag, a Samsung Galaxy S6, and a Samsung Galaxy Gear, this dataset spans multiple zero-interaction security use cases in an office environment.
These four datasets provide a variety of sensor types, user activities, and sampling rates, allowing us to explore how intrinsic biases and correlations manifest across different real-world scenarios. Next, we detail how we preprocess and aggregate this data to form global distributions for our entropy analyses.
4 Entropy Analysis
In this section, we analyze the intrinsic entropy of sensor data under the threat model described in §3. We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our single-sensor findings. Later sections examine multi-sensor joint entropies and the complexities of sensor fusion.
4.1 Pre-processing
A crucial, yet underexplored, issue in prior work (e.g. [19, 25, 26, 31]) is how to convert inherently continuous sensor outputs into suitable discrete values for entropy estimation. For example, Shannon entropy and min-entropy, as defined in Eqs. 3 and 5, rely on discrete random variables. Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution. However, a sensor’s advertised resolution (e.g. 12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform or truly high-fidelity coverage across its range. Everyday usage introduces biases and clustering, resulting in some measurements occurring far more frequently than others. For instance, UCI-HAR data shows accelerometer readings widely spread in certain axes, while approximately 60% of gyroscope readings hover near zero (see Figure 2). Such heavy skew or near-constant values radically diminish real-world entropy, which would otherwise benefit from uniformly distributed values.
(a) Acc. $x$ axis.
(b) Acc. $y$ axis.
(c) Acc. $z$ axis.
(d) Gyro. $x$ axis.
(e) Gyro. $y$ axis.
(f) Gyro. $z$ axis.
(g) Accelerometer.
(h) Gyroscope.
(i) Light.
(j) Linear Acceleration.
(k) Magnetometer.
(l) Rotation Vector.
Figure 1: Global sensor data CDFs – UCI-HAR (a–f) and Relay (g–l) datasets.
[…]
$h=2\cdot\frac{IQR(x)}{n^{1/3}}$ (8)
4.2 Single Sensors
Given the biases discussed above, it is unsurprising that some sensor readings exhibit high predictability in the worst-case sense. We illustrate this by calculating individual-sensor entropy across multiple datasets, splitting multi-dimensional readings (e.g., triaxial accelerometer or gyroscope) into separate axes following Voris et al. [19]. The results are given in Table 1. We note that in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors. Rather, the authors preprocess triaxial data into its vector magnitude, i.e. $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$. We give this as “X.Mag” for a given sensor X. For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this abstract sensor modality.
Several clear patterns emerge from Table 1. Some sensors (e.g. certain accelerometer axes in SHL or PerilZIS) exhibit moderate min-entropies (4–6 bits). Other sensors, particularly gyroscope axes, (see UCI-HAR), show values below 3 bits, indicating high predictability in their most frequent readings. Shannon entropy values ($H_{1}$) can be fairly high (up to 10 bits in some cases), whereas min-entropy ($H_{\infty}$) is often much lower. This gap reflects distributions where a few outcomes dominate, thereby driving worst-case unpredictability down even if the average-case picture is more favorable. Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own. In the next section, we examine whether combining multiple modalities can meaningfully increase this worst-case unpredictability or whether correlated biases persist across different sensor streams.
Dataset
UCI-HAR SHL Relay* PerilZIS
Sensor $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
Acc.X 8.488 7.080 5.876 3.729 11.557 8.732 7.487 4.543 13.012 9.292 6.359 3.626
Acc.Y 8.243 7.231 6.847 5.694 11.425 8.928 7.717 4.500 9.549 5.873 4.483 2.889
Acc.Z 8.455 7.397 7.069 6.020 10.428 7.627 6.366 3.785 9.817 6.671 5.403 4.002
Acc.Mag 8.895 6.284 4.819 3.489 14.583 10.136 8.710 6.435 10.145 6.843 5.808 4.538 13.328 8.273 7.115 4.526
Gyro.X 8.683 5.430 3.504 1.929 15.024 10.532 8.107 4.993 14.528 7.231 4.454 2.805
Gyro.Y 8.439 5.023 3.461 2.300 15.085 10.283 7.601 4.827 14.078 6.715 4.039 2.529
Gyro.Z 8.714 5.675 3.948 2.363 15.281 10.070 5.708 3.083 13.961 6.463 3.836 2.434
Gyro.Mag 8.414 5.759 4.130 2.537 12.123 7.816 5.728 3.699 7.954 4.751 3.442 2.083 14.166 5.565 1.932 0.969
Mag.X 12.845 8.840 8.386 6.374 10.767 7.639 6.816 4.883
Mag.Y 12.263 8.737 8.314 6.223 10.179 7.622 6.605 4.405
Mag.Z 12.516 8.586 8.217 6.228 10.129 7.507 6.726 4.448
Mag.Mag 13.558 9.436 8.771 7.148 7.972 6.147 5.617 4.254 10.293 7.329 6.489 4.454
Rot. Vec. 8.725 7.721 5.970 3.220 5.000 3.307 1.965 1.021
Grav.X 9.014 8.482 7.266 4.299
Grav.Y 9.338 8.770 7.602 4.453
Grav.Z 8.180 7.193 5.418 3.036
Grav.Mag 14.373 7.988 7.227 6.242 7.794 6.325 5.864 4.532
LinAcc.X 15.260 10.077 7.621 5.116
LinAcc.Y 14.859 10.116 7.639 5.224
LinAcc.Z 14.377 9.951 7.605 4.543
LinAcc.Mag 12.777 7.968 5.752 3.420 9.175 6.385 5.424 4.222
Light 7.200 5.331 4.507 3.206 12.152 7.940 7.137 4.552
Humidity 7.943 7.048 6.774 5.546
Temp. 7.295 4.753 2.611 1.332 8.484 7.416 6.941 5.449
Pressure 9.461 8.170 7.723 6.237 8.044 7.006 6.370 5.073
Mean 8.541 6.235 4.957 3.508 13.188 9.266 7.178 4.483 7.891 5.584 4.661 3.408 11.277 7.224 5.717 3.912
S.D. 0.207 0.904 1.461 1.574 1.993 1.148 1.115 1.018 1.612 1.227 1.474 1.379 2.312 0.900 1.526 1.266
  • *: The authors preprocess triaxial data into its vector magnitude.
Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data for that modality and dataset.
4.3 Multi-modal Sensors
Many sensor-based security proposals [6, 5, 4, 2, 20] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one. A naïve model assumes each sensor is statistically independent, yielding the simplified relation $H(X_{1},X_{2},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, real-world sensor outputs often exhibit significant interdependence. For example, linear acceleration, rotation vector, and gravity in the Android ecosystem are often derived from the same underlying motion sensors, causing strong correlations. Figure 2 illustrates the correlation matrices of the sensors we examine across our four datasets, revealing that many modalities are far from independent.
(a) UCI-HAR
(b) PerilZIS
(c) Relay
(d) SHL
Figure 2: Sensor correlation matrices for each dataset.
4.3.1 Complexity challenges.
Unfortunately, directly estimating the joint entropy of discrete sensor data often becomes computationally infeasible as the number of modalities increases, due to the curse of dimensionality. Specifically, for $n$ modalities, where each modality $X_{i}$ is discretised into $b_{i}$ bins (with $\{b_{1},b_{2},\dots,b_{n}\}\subseteq B$), the total number of joint bins scales as $\prod^{n}_{i=1}b_{i}$. Applying Freedman–Diaconis binning rules typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially. Consequently, computing a full joint entropy estimation, as in Eq. 6, becomes intractable. Another concern is the combinatorial explosion of sensor combinations. Let us assume that one has 19 modalities, as in Table 1, then the total number of combinations to analyse is equal to $202^{19}-$ 20, skipping single sensors and the empty set. For even moderately large datasets, iterating over every sensor subset and computing its full joint entropy is prohibitively expensive. Addressing these problems is challenging. One mitigation is to reduce the number of bins per modality. However, this inevitably risks oversimplifies the distribution and underestimating its joint entropy. Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset. Further scaling became infeasible. We also observed that imposing these smaller bin counts (rather than the thousands implied by Freedman–Diaconis) reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1.
4.3.2 Chow-Liu approximation.
To address the problem of high-dimensional entropy estimation, we adopt the Chow-Liu algorithm [44] to approximate the joint probability distribution of multiple sensors using a tree-structured Bayesian network. Specifically, Chow-Liu constructs a maximum-weight spanning tree $\pi$ over the $n$ modalities, where the edge weights correspond to the mutual information between pairs of variables. This tree structure minimises the Kullback-Leibler divergence (Def. 2.3) between the true joint distribution and the tree-structured approximation. The joint probability is then estimated as:
$p_{\pi}(x_{1},\dots,x_{n})=p(x_{r})\prod_{i\neq r}p(x_{i}\;|\;x_{\pi(x)})$ (9)
Where $\pi(i)$ denotes the parent of $X_{i}$ in the tree, and $r$ is the tree’s root node. This model efficiently captures the dominant pairwise dependencies, which is particularly beneficial when exploring the combinatorial relationships of different modalities. The use of Chow-Liu trees was also suggested by Buller and Kaufer [35] for estimating the entropy of multivariate data sources where the range of possible values is high. In our Python implementation, we use the pgmpy [45] library’s TreeSearch module. With the inferred Bayesian network in hand, we can compute the max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$) and min-entropies ($H_{\infty}$) without resorting to summation over the full exponential state space. Chow–Liu’s acyclic, singly connected structure is crucial: each node has at most one parent, so one can traverse the tree top-down (from root) or bottom-up to accumulate either sums of probabilities (e.g. for collision entropy) or maximum probabilities (e.g. min-entropy). This significantly reduces computation time compared to naïve enumeration.
Our entropy analysis framework systematically evaluates the joint entropy over all valid combinations of sensor modalities. The powerset of the sensor set is generated (with restrictions on minimum and maximum combination lengths) and processed in parallel using Python’s multiprocessing module. Each subset is independently analysed: quantising the data, inferring a Chow–Liu tree, and computing the suite of entropy metrics. Processing all four datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
4.3.3 Results.
The results are given in Tables 2 to 5, listing the top 10 sensor combinations for each dataset in descending order of min-entropy. The full sensor set consistently achieves the highest entropy values. This behavior is anticipated: adding more modalities increases the theoretical upper bound on entropy (e.g., $H_{0}$), and it does not lower the min-entropy. We observe that removing certain modalities often has a negligible impact on worst-case unpredictability. For example, in the SHL dataset (Table 5), while the full sensor suite attains $H_{0}=158.601$ bits and $H_{1}=82.301$ bits, the min-entropy remains at a constant $H_{\infty}=21.289$ bits across both the full sensor set and the top-performing subsets. This consistency indicates that, although additional sensors expand the total support (reflected in higher $H_{0}$), they do not necessarily enhance worst-case unpredictability ($H_{\infty}$). Similarly, in Table 3 for the Relay dataset, the combination (Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits. Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$.These observations suggest that carefully chosen sensor subsets can yield nearly optimal min-entropy while reducing the number of modalities and associated overhead (e.g. battery consumption). We provide the full sensor combination results in our open-source repository.
Table 2: Top 10 best-performing sensor combinations (UCI-HAR; in bits).
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 48.061 25.089 17.116 11.008
(Acc.{x,y,z}, Gyro.{y,z}) 39.403 22.835 15.999 10.113
(Acc.{x,y,z}, Gyro.{x,y}) 39.886 21.456 15.332 9.900
(Acc.{x,y,z}, Gyro.{x,z}) 40.222 21.470 15.040 9.411
(Acc.{y,z}, Gyro.{x,y,z}) 40.228 21.306 14.698 9.274
(Acc.{x,z}, Gyro.{x,y,z}) 40.293 21.260 14.575 9.154
(Acc.{x,y,z}, Gyro.y) 31.228 18.804 13.893 9.065
(Acc.{x,y}, Gyro.{x,y,z}) 40.273 21.155 14.208 8.721
(Acc.{x,y,z}, Gyro.z) 31.564 18.775 13.833 8.682
(Acc.{y,z}, Gyro.{y,z}) 31.570 18.583 13.400 8.386
Table 3: Top 10 best-performing sensor combinations (Relay; in bits).
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 45.794 25.036 15.725 8.092
(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 44.795 24.963 15.334 7.859
(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 38.385 21.590 14.783 7.702
(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 37.026 21.001 14.553 7.702
(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 36.092 20.794 14.450 7.673
(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 37.385 21.517 14.428 7.469
(Acc., Gyro., Light, Mag., Rot. Vec.) 36.026 20.924 14.270 7.465
(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 35.092 20.720 14.147 7.439
(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 41.094 22.794 14.437 7.338
(Acc., Grav., Light, Mag., Rot. Vec.) 29.617 17.555 13.244 7.312
Table 4: Top 10 best-performing sensor combinations (PerilZIS; in bits).
Modalities $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 83.662 36.998 33.682 23.926
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 76.907 34.258 31.510 23.835
(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.246 31.509 23.835
(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.820 34.243 31.508 23.829
(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.229 31.501 22.945
(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 71.384 32.586 30.190 22.945
(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 68.568 34.218 31.500 22.945
(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 68.650 34.215 31.499 22.945
(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 67.214 32.575 30.189 22.945
(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 76.618 33.816 31.035 21.915
Table 5: Top 10 best-performing sensor combinations (SHL; in bits).
Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
All sensors 158.601 82.301 39.320 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 148.995 78.624 39.276 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 148.759 78.477 39.272 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 148.587 78.380 39.249 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 148.952 78.135 39.235 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 139.153 74.800 39.175 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 138.981 74.703 39.128 21.289
(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 138.745 74.556 39.117 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 139.346 74.458 39.100 21.289
(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 139.109 74.311 39.087 21.289
5 Evaluation
This section critically evaluates our sensor entropy findings using existing literature and relevant standards. We then discuss potential mitigations such as randomness extractors, outline inherent limitations of our study and sensor-based approaches, and propose avenues for future enhancements.
5.1 Comparison with Related Work
Prior studies have claimed that mobile sensors can serve as a robust source of security or randomness. Much work has relied on machine-learning-based metrics as a proxy to evaluate such claims [6, 7, 4, 2], with a smaller subset using more established entropy metrics. Li et al.[21] report 32–38 bits of Shannon entropy for some motion-based secure pairing schemes, and Wu et al.[22] estimate 51–54 bits. However, dedicated entropy analyses frequently demonstrate that individual sensors confer very little min-entropy, especially under worst-case assumptions [25, 26, 19, 3]. Our results, derived from a systematic analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
Single sensors yield low entropy. In line with prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits). Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g. some single-sensor readings yield only 1–3 bits of min-entropy). This highlights a major gap between average-case unpredictability and worst-case security, which is an important distinction that many earlier evaluations do not fully capture.
Multiple sensors improve worst-case entropy, but not dramatically. While combining more modalities typically raises the upper bound (e.g. Hartley or Shannon entropy), our experiments reveal that min-entropy gains are far smaller than one might hope. In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution.
Overall, our findings place stricter bounds on the real benefits of mobile-sensor-based approaches. While sensor data can be effective for classification-based tasks such as continuous authentication, where strong assurances are not necessarily required, or simple proximity checks, the suitability of sensor data as an attack-resistant random source remains doubtful unless carefully engineered and rigorously evaluated. Min-entropy is critical in scenarios where an adversary can guess the most likely states first; our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further. Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone replacement for dedicated, cryptographically validated entropy sources.
5.1.1 Can Extractors Salvage Weak Sources?
One might hope that cryptographic extractors (e.g. Von Neumann extractors or more advanced schemes [46, 47, 48]) could mitigate low-entropy sensor data. Indeed, extractors are designed to reduce bias in a noisy or skewed source. However, they cannot increase the total amount of randomness beyond the source’s intrinsic min-entropy. In other words, if the combined sensor data only provides around 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more. From a security standpoint, this means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance). Thus, while extraction can help if the original data has enough min-entropy to meet system requirements, it cannot elevate a weak source—one that fundamentally lacks sufficient min-entropy—to the higher security thresholds often demanded by rigorous standards.
5.2 Limitations
Despite our comprehensive multivariate analysis, this work has some limitations. Firstly, although we analyse four large datasets, they do not fully capture contexts all possible contexts. For example, work has suggested that performing dedicated movements (e.g. gestures) can increase the amount of usable entropy from motion sensors, on the order of magnitude of 5–6 bits [25, 19]. Our datasets do not cover such dedicated movements; it is possible that the reported resuts are an underestimation of entropy for motion-based sensors. Secondly, our choice of Freedman-Diaconis binning and Chow-Liu trees is a pragmatic compromise. Smaller bin sizes tend to over-simplify the distribution and underestimate entropy, whereas larger bins can lead to computational blowup. Although our approach outperforms naive joint entropy estimation, it is still an approximation. Thirdly, we largely focus on a data-centric view of entropy and do not deeply explore sophisticated adversarial strategies, such as sensor spoofing [20], fault injection attacks [34] or cross-modality correlation attacks (e.g. see [49] for such attacks against wireless body area networks). These threats may significantly reduce the usable entropy of sensors even further.
6 Conclusion
This paper provides a comprehensive analysis of sensor-derived entropy across multiple datasets and modalities. Our results expose a tension between the perceived and actual strength of sensor data for security applications. Even in the best-performing sensor combinations, seemingly suitable results using one metric collapse to insecure levels when using standard worst-case metrics. Notably, modalities that yield ‘good’ max- or Shannon entropies, representing the best- and average-case unpredictability, have insecure worst-case min-entropies. Consequently, sensor modalities that may appear robust have biases that may enable adversaries to predict the most probable values with minimal effort.
Our findings also challenge the prevailing evaluation approach that machine learning metrics (e.g., accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals. The vulnerability of mobile sensors to biased distributions significantly undermine their effectiveness as reliable entropy sources. We also cast doubt on the use of sensors with respect to their non-stationary and lack of reproducibility. These issues collectively contradict the criteria articulated in frameworks such as NIST SP 800-90B, which emphasise noise-source stationarity and protection from external influence. The effectiveness of countermeasures remains an open research challenge. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, and ultimately pave the way for safer and more robust designs in mobile-device security.
In future work, we consider that a dynamic analysis is important to assess the stationarity issues with sensor data, where entropy varies according to user behavior or environment. Moreover, a user study could yield empirical bounds on how finely humans can intentionally manipulate motion sensors, revealing more realistic limits to sensor-based randomness in real-world scenarios. For example, our binning decision was a statistical one, rather one that reflects human usage; it is possible that real-world usage may reduce the resolution of useful sensor data, thus reducing entropy. Overall, while sensor-based data can reach limited levels of entropy under favourable conditions, the road to making such sources systematically secure and robust is long. The key takeaway is that researchers and engineers should resist the temptation to treat sensor readings as bona fide random bits. Substantial work is needed before sensors can be considered as appropriate sources of randomness for security-critical applications.
6.0.1 Acknowledgments
Carlton Shepherd has received funding from the UK EPSRC ‘Chameleon’ project (EP/Y030168/1). The authors would like to thank Darren Hurley-Smith for insightful conversations on the topic.

### v3

@@
 A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference.
-Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while newer techniques incorporate ambient phenomena—characteristics of magnetic fields [2] and thermal fluctuations [3]—to mitigate relay attacks or reduce user-driven authentication prompts.
+Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] or reduce user-driven authentication prompts [7, 8, 9, 10, 11].
 Despite these advancements, fundamental questions remain.
@@
 Many existing studies rely on heuristic assessments or machine learning classifiers, e.g.
-[4, 6, 7, 2, 5], that do not address core cryptographic requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
+[4, 13, 14, 2, 12], that do not address core security requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
 Our analysis reveals systemic limitations.
@@
 Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] have been used to evaluate the effectiveness of distinguishing between legitimate and illegitimate samples.
-User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [11, 1].
-However, machine learning-based evaluation metrics do not measure intrinsic entropy—a crucial factor in determining whether sensor data is a suitable source of randomness for security applications.
+(User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].) We observe that machine learning-based evaluation metrics do not directly measure entropy, which is a crucial factor in determining whether sensor data is a suitable source of randomness for security applications.
 Some work has attempted to assess the entropy of sensor signals, but this represents a minority.
@@
 We aim to determine the global (i.e., population-level) entropy characteristics of various mobile sensors under ordinary usage conditions, rather than focusing on per-user or scenario-specific differences.
-This choice reflects common real-world deployments, which often must accommodate a wide range of behaviors and environments with minimal tailoring.
+This choice reflects common real-world deployments, which must accommodate a wide range of behaviors and environments.
 Our approach involves five main stages:
@@
 Our entropy analysis framework systematically evaluates the joint entropy over all valid combinations of sensor modalities.
-The powerset of the sensor set is generated (with restrictions on minimum and maximum combination lengths) and processed in parallel using Python’s multiprocessing module.
+The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module.
 Each subset is independently analysed: quantising the data, inferring a Chow–Liu tree, and computing the suite of entropy metrics.

### v4

@@
 A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference.
-Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] or reduce user-driven authentication prompts [7, 8, 9, 10, 11].
-Despite these advancements, fundamental questions remain.
-Multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated.
+Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11].
+Despite these advancements, fundamental issues remain: multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated.
 Many existing studies rely on heuristic assessments or machine learning classifiers, e.g.
-[4, 13, 14, 2, 12], that do not address core security requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
+[4, 13, 14, 2, 12], that do not address critical security questions.
+That is, firstly, how much entropy do sensors truly provide?
+And, secondly, to what extent do multi-modal sensor combinations provide security gains?
 Our analysis reveals systemic limitations.
-Commodity sensors exhibit significant biases, with gyroscopes yielding min-entropy as low as 0.96 bits—well below cryptographic thresholds.
-Although multi-modal sensor fusion mitigates some of these issues, non-uniform distributions and inter-sensor correlations reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy.
-These findings challenge the notion that increasing the number of sensors automatically strengthens security, and underscore the inadequacy of raw sensor data as a reliable entropy source.
-Our contributions are:
+Commodity sensors exhibit significant biases, with gyroscopes yielding min-entropy as low as 0.96 bits.
+While multi-modal sensor usage somewhat addresses these issues, non-uniform distributions and inter-sensor correlations reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy.
+These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using raw sensor data as a dependable entropy source.
+Our contributions are as follows:
   • We introduce the first systematic approach to evaluating sensor entropy across diverse modalities and datasets using various entropy metrics (max, Shannon, collision, and min-entropy).
@@
   • Ultimately, we advise against relying on commodity sensors as primary sources of randomness for security-critical applications, both on a single- and multiple-sensor basis.
-The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics.
+The rest of this paper is organised as such: §2 discusses sensor-based security mechanisms and established entropy metrics.
 §3 explains our experiment design for entropy estimation, including the threat model and dataset selection.
@@
 We conclude in §6 with recommendations for further work.
-Our analysis framework is made publicly available to foster future research.
+Our analysis work is released publicly to foster future research.
 [^1] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@
 2.1 Mobile Sensor-based Security Applications
-One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [15, 13, 14, 4].
+One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [15, 13, 14, 4, 2].
 Mehrnezhad et al.[13] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together.
@@
 [14] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device.
-Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [16, 17, 18, 8, 9, 19, 10, 11], where sensor data is used to authenticate users passively without explicit interaction.
-(with up to 99% classification accuracy in some instances [17]).
-Mobile sensors have also been proposed as enabling novel device pairing schemes for secure communication, e.g.
-[1, 20].
-The notion of ‘hardness’ is typically inferred based on a machine learning-based analysis; it is not grounded in standard quantitative methods used in the field of authentication, such as Shannon [21] and min-entropy [22].
-A general approach to follows one of those given by Shrestha et al.
-[3].
-Here, sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g.
-cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13].
-Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] have been used to evaluate the effectiveness of distinguishing between legitimate and illegitimate samples.
-(User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].) We observe that machine learning-based evaluation metrics do not directly measure entropy, which is a crucial factor in determining whether sensor data is a suitable source of randomness for security applications.
+Shrestha et al.
+[2] explore four environmental modalities, including temperature and humidity, for proximity detection.
+In this work, single sensors yields 0.733–0.881 F1-score, while combining multiple increases the overall performance to 0.913–0.957.
+Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [16, 17, 18, 8, 9, 19, 10, 11], where sensor data is used to authenticate users passively without explicit interaction (with up to 99% accuracy claimed in some work [17]).
+Sensors have also been used to underpin the security of novel device pairing schemes, whether as a primary [20] or second line of authentication [1].
+These schemes generally rely on detecting similar motion patterns between two devices using joint accelerometer and gyroscope sensor measurements in order to provide evidence of co-location.
+The notion of ‘hardness’ is typically inferred using model evaluation metrics; it is not grounded in standard methods used in the field of authentication, such as Shannon [21] and min-entropy [22].
+A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g.
+cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2].
+Features are subsequently classified using simple threshold-based or supervised classification models, e.g.
+Support Vector Machine (SVM) and Random Forest.
+Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are then used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples.
+(User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].)
+We observe that the vast majority of proposals do not directly measure entropy, which is a crucial factor in determining whether sensor data cannot be easily predicted by an adversary.
 Some work has attempted to assess the entropy of sensor signals, but this represents a minority.
 T2Pair by Li et al.
-[23] 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al.
+[23] yielded 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al.
 [24] reporting 51–54 bits of Shannon entropy.
 Even in the best cases, this is low relative to modern cryptographic standards.
-We point to adjacent work that has questioned the utility of mobile sensors.
+[^2] See AIS 20/31 [25]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator.
+Lastly, we point to adjacent work that has questioned the utility of mobile sensors.
 Markantonakis et al.
@@
 [5] and Shepherd et al.
-[6], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints.
-Using consumer off-the-shelf devices, the work reports that sensors available through the Android platform could not be used to reliably detect man-in-the-middle attacks in NFC contactless transactions within the 500ms time window required by the EMV payment protocol.
-Under those constraints, the effectiveness drastically decreased compared to previously reported results.
-Indeed, it was lower to the extent such that sensors could not be recommended for mobile transactions without posing significant usability and security issues in practice.
+[6], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints for NFC-based contactless transactions.
+Using a 500ms window for data acquisition, as stipulated by the EMV payment protocol, 0.179–0.246 EER was reported depending on the given sensor combination.
+Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
 2.2 Sensor Entropy Analyses
@@
 [22] investigated the use of accelerometers as true random number generators (TRNGs), specifically aboard a WISP RFID tag and a Nokia N97 mobile phone.
-The authors find that min-entropy—defined in §2.3—that min-entropy is proportional to the motion applied to the device, and that stationary movement has the lowest min-entropy.
+The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, and that stationary movement has the lowest min-entropy.
 Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC) are considered other significant influences on the generated entropy.
@@
 Mobile sensors depend heavily on user behaviour and environmental, introducing biases and correlations absent in controlled entropy sources.
-Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic assessments or those based on machine learning-based evaluation metrics [8, 11].
+Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic or model evaluation metrics [8, 11].
 Such approaches do not account well for skewed distributions and other biases in the underlying data.
@@
 We draw attention to four special cases of $\alpha$ that are widely used in the literature.
-Firstly, the Hartley (Max) Entropy ($\alpha=0$), given in Eq.
+Firstly, the Hartley (Max) Entropy, given in Eq.
 2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
-$H_{0}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
+$H_{0}(X)\equiv\lim_{\alpha\to 0}H_{\alpha}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
 Second is the Shannon Entropy (Eq. 3), which corresponds to the classical definition of entropy in information theory, and is the limit of $H_{\alpha}$ as $\alpha\to 1$.
-$H_{1}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
+$H_{1}(X)\equiv\lim_{\alpha\to 1}H_{\alpha}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
 Another case is the Collision Entropy ($\alpha=2$), quantifying the probability of “collisions” of multiple draws from $X$. This is given in Eq. 4.
@@
 5), i.e.
-the value at the limit $\alpha\to\infty$.
-$H_{\infty}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
+the value in the limit $\alpha\to\infty$.
+$H_{\infty}(X)\equiv\lim_{\alpha\to\infty}H_{\alpha}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
 Note that $H_{\alpha}$ is a non-increasing function of $\alpha$, i.e.
@@
 We consider an adversary whose primary objective is to compromise sensor-based key generation or authentication schemes by exploiting the inherent biases or predictable patterns within sensor data.
-In particular, we assume the adversary has full or near-full knowledge of the probability distribution of sensor outputs.
+We assume the adversary has full or near-full knowledge of the probability distribution of sensor outputs.
 They may obtain this through statistical modeling of publicly available datasets, prolonged observation of target users, or eavesdropping on sensor data streams.
-We exclude adversarial capabilities such as fault-injection attacks or hardware attacks, e.g.
+We exclude capabilities such as fault-injection and other hardware attacks, e.g.
 see [34].
-While such attacks could further reduce the effective entropy space, we regard them as a higher level of sophistication—an important future research direction but not the focus here.
-Lastly, rather than brute-forcing all possible sensor outputs, we assume a more strategic approach: the adversary uses its knowledge of the sensor’s distribution to prioritise the most probable (or most easily predicted) outcomes.
+While such attacks could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard this out-of-scope for this work.
+Furthermore, rather than brute-forcing all possible sensor outputs, we assume a more strategic approach.
+The adversary uses its knowledge of the sensor’s distribution to prioritise the most probable outcomes.
 By iterating guesses in order of descending probability, they minimise their expected effort.
-This threat model represents a realistic adversary who can capitalise on statistical biases in sensor data but does not directly sabotage or physically compromise the device.
-Our goal is to evaluate whether typical sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist such knowledge-driven attacks.
+This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically.
+Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist knowledge-driven attacks.
 3.2 High-level Methodology
@@
 For sensors that are inherently discrete or quantised (e.g., integer output ranges), we simply count occurrences.
-For sensors that produce (quasi-)continuous values, we rely on quantisation strategies (Freedman–Diaconis binning) to partition the output space and approximate an empirical probability mass function.
+For sensors that produce (quasi-)continuous values, we rely on quantisation using Freedman–Diaconis binning to partition the output space and approximate an empirical probability mass function.
   3. 3.
@@
 Data includes triaxial accelerometer and gyroscope signals.
-  • University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz.
-Multiple Huawei Mate 9 phones attached across the users’ body in a body area network (BAN).
-The publicly available SHL Preview dataset is used, comprising three recording-days per user (a total of 59 hours of data).
-To scope this study, we use the motion sensors dataset from the handheld mobile phone as a good fit with related work (see §2).
+  • University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz from an Huawei Mate 9 smartphone.
+The publicly available SHL Preview dataset is used, comprising three recording-days per user (59 hours of data in total).
+To scope this study, we use the dataset from the handheld mobile phone as a good fit with related work (see §2).
   • Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés).
@@
 4 Entropy Analysis
-In this section, we analyze the intrinsic entropy of sensor data under the threat model described in §3.
-We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our single-sensor findings.
-Later sections examine multi-sensor joint entropies and the complexities of sensor fusion.
+In this section, we analyse the intrinsic entropy of sensor data under the threat model described in §3.
+We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our findings for our single- and multi-sensor analyses.
 4.1 Pre-processing
@@
 [22, 26, 27, 32]) is how to convert inherently continuous sensor outputs into suitable discrete values for entropy estimation.
-For example, Shannon entropy and min-entropy, as defined in Eqs.
+For example, Shannon and min-entropy, as defined in Eqs.
 3 and 5, rely on discrete random variables.
 Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution.
-However, a sensor’s advertised resolution (e.g.
-12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform or truly high-fidelity coverage across its range.
+Yet, a sensor’s advertised resolution (e.g.
+12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform coverage across its range.
 Everyday usage introduces biases and clustering, resulting in some measurements occurring far more frequently than others.
-For instance, UCI-HAR data shows accelerometer readings widely spread in certain axes, while approximately 60% of gyroscope readings hover near zero (see Figure 2).
-Such heavy skew or near-constant values radically diminish real-world entropy, which would otherwise benefit from uniformly distributed values.
+For instance, UCI-HAR data shows accelerometer readings concentrated in certain areas, and approximately 60% of gyroscope readings hover near zero (see Figure 2).
+Such skew and bias radically diminishes entropy compared to uniformly distributed values.
 (a) Acc. $x$ axis.
@@
 4.2 Single Sensors
-Given the biases discussed above, it is unsurprising that some sensor readings exhibit high predictability in the worst-case sense.
-We illustrate this by calculating individual-sensor entropy across multiple datasets, splitting multi-dimensional readings (e.g., triaxial accelerometer or gyroscope) into separate axes following Voris et al.
+Given the biases discussed above, it is inevitable that some sensor readings will exhibit relatively high predictability.
+To quantify this, we calculate individual-sensor entropies across multiple datasets.
+The results are given in Table 1.
+For multi-dimensional modalities (e.g.
+triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al.
 [22].
-The results are given in Table 1.
-We note that in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors.
-Rather, the authors preprocess triaxial data into its vector magnitude, i.e.
+We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors.
+Rather, the authors have already preprocessed triaxial data into its vector magnitudes, i.e.
 $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$.
 We give this as “X.Mag” for a given sensor X.
-For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this abstract sensor modality.
+For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this new synthetic modality.
 Several clear patterns emerge from Table 1.
@@
   • *: The authors preprocess triaxial data into its vector magnitude.
-Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data for that modality and dataset.
+Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data.
 4.3 Multi-modal Sensors
-Many sensor-based security proposals [13, 12, 4, 2, 3] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one.
-A naïve model assumes each sensor is statistically independent, yielding the simplified relation $H(X_{1},X_{2},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
+Several sensor-based security proposals [13, 12, 4, 2, 3] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one.
+This section will examine that claim.
+A naïve approach to analysing the entropy of multiple sensors is to first assume that each sensor is statistically independent, yielding the simplified relation $H(X_{1},X_{2},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
 However, real-world sensor outputs often exhibit significant interdependence.
-For example, linear acceleration, rotation vector, and gravity in the Android ecosystem are often derived from the same underlying motion sensors, causing strong correlations.
-Figure 2 illustrates the correlation matrices of the sensors we examine across our four datasets, revealing that many modalities are far from independent.
+For example, the linear acceleration, rotation vector, and gravity sensors in the Android ecosystem can be implemented in software using data from accelerometers and gyroscopes, creating a source of correlation [44].
+Figure 2 shows the correlation matrices of the modalities across our four datasets, revealing that many modalities are far from independent.
 (a) UCI-HAR
@@
 4.3.1 Complexity challenges.
-Unfortunately, directly estimating the joint entropy of discrete sensor data often becomes computationally infeasible as the number of modalities increases, due to the curse of dimensionality.
+Unfortunately, directly estimating the joint entropy of discrete sensor data becomes computationally infeasible as the number of modalities increases due to the curse of dimensionality.
 Specifically, for $n$ modalities, where each modality $X_{i}$ is discretised into $b_{i}$ bins (with $\{b_{1},b_{2},\dots,b_{n}\}\subseteq B$), the total number of joint bins scales as $\prod^{n}_{i=1}b_{i}$.
@@
 One mitigation is to reduce the number of bins per modality.
-However, this inevitably risks oversimplifies the distribution and underestimating its joint entropy.
+However, this inevitably risks oversimplifying the distribution and underestimating its joint entropy.
 Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset.
@@
 In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution.
-Overall, our findings place stricter bounds on the real benefits of mobile-sensor-based approaches.
-While sensor data can be effective for classification-based tasks such as continuous authentication, where strong assurances are not necessarily required, or simple proximity checks, the suitability of sensor data as an attack-resistant random source remains doubtful unless carefully engineered and rigorously evaluated.
-Min-entropy is critical in scenarios where an adversary can guess the most likely states first; our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further.
-Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone replacement for dedicated, cryptographically validated entropy sources.
+Min-entropy is critical in scenarios where an adversary can guess the most likely states first.
+Overall, our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches.
+We estimate that even the most complicated multi-modal combination provide relatively little entropy.
+It is important to note that our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further.
+Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone source.
+They might be useful for tasks where strong assurances are not required, e.g.
+simple proximity checks; however, the suitability of sensor data as a robust source of unpredictability remains highly doubtful.
 5.1.1 Can Extractors Salvage Weak Sources?
@@
 Von Neumann extractors or more advanced schemes [47, 48, 49]) could mitigate low-entropy sensor data.
-Indeed, extractors are designed to reduce bias in a noisy or skewed source.
-However, they cannot increase the total amount of randomness beyond the source’s intrinsic min-entropy.
-In other words, if the combined sensor data only provides around 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more.
+Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of randomness beyond the source’s intrinsic min-entropy.
+That is, if the combined sensor data only provides around 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more.
 From a security standpoint, this means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance).
@@
 5.2 Limitations
-Despite our comprehensive multivariate analysis, this work has some limitations.
-Firstly, although we analyse four large datasets, they do not fully capture contexts all possible contexts.
-For example, work has suggested that performing dedicated movements (e.g.
-gestures) can increase the amount of usable entropy from motion sensors, on the order of magnitude of 5–6 bits [25, 22].
-Our datasets do not cover such dedicated movements; it is possible that the reported resuts are an underestimation of entropy for motion-based sensors.
+Despite our comprehensive analysis, this work has some limitations.
+Firstly, while we analyse four large datasets, they do not fully capture contexts all possible contexts.
+Some work has suggested that performing dedicated movements (e.g.
+gestures) can increase the amount of usable entropy from motion sensors by 5–6 bits [26, 22].
+Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action.
 Secondly, our choice of Freedman-Diaconis binning and Chow-Liu trees is a pragmatic compromise.
 Smaller bin sizes tend to over-simplify the distribution and underestimate entropy, whereas larger bins can lead to computational blowup.
-Although our approach outperforms naive joint entropy estimation, it is still an approximation.
-Thirdly, we largely focus on a data-centric view of entropy and do not deeply explore sophisticated adversarial strategies, such as sensor spoofing [3], fault injection attacks [34] or cross-modality correlation attacks (e.g.
-see [49] for such attacks against wireless body area networks).
-These threats may significantly reduce the usable entropy of sensors even further.
+As such, although our approach outperforms naïve joint entropy estimation, it is still an approximation.
+Thirdly, we largely focus on a data-centric view of entropy and do not examine other threats, such as sensor spoofing [3], hardware attacks [34] and cross-device correlation attacks (see [50] against wireless body area networks).
+It is possible that these threats may reduce the entropy of sensors even further.
 6 Conclusion
@@
 Consequently, sensor modalities that may appear robust have biases that may enable adversaries to predict the most probable values with minimal effort.
-Our findings also challenge the prevailing evaluation approach that machine learning metrics (e.g., accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals.
+Our findings also challenge the prevailing orthodoxy that model evaluation metrics (e.g.
+accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals.
 The vulnerability of mobile sensors to biased distributions significantly undermine their effectiveness as reliable entropy sources.

### v5

@@
   • We empirically demonstrate how inter-sensor correlations and biases erode entropy, casting doubt on the proposition that using multiple sensors adds substantially to security.
-  • Ultimately, we advise against relying on commodity sensors as primary sources of randomness for security-critical applications, both on a single- and multiple-sensor basis.
-The rest of this paper is organised as such: §2 discusses sensor-based security mechanisms and established entropy metrics.
+  • Ultimately, we advise against relying on commodity sensors as sources of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
+The rest of this paper is organised in the following way: §2 discusses sensor-based security mechanisms and established entropy metrics.
 §3 explains our experiment design for entropy estimation, including the threat model and dataset selection.
@@
 These schemes generally rely on detecting similar motion patterns between two devices using joint accelerometer and gyroscope sensor measurements in order to provide evidence of co-location.
-The notion of ‘hardness’ is typically inferred using model evaluation metrics; it is not grounded in standard methods used in the field of authentication, such as Shannon [21] and min-entropy [22].
+The notion of ‘hardness’ is typically inferred through model evaluation metrics.
 A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g.
 cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2].
-Features are subsequently classified using simple threshold-based or supervised classification models, e.g.
-Support Vector Machine (SVM) and Random Forest.
-Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are then used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples.
+Features are then classified using simple threshold-based or supervised classification models, e.g.
+Support Vector Machines (SVM) and Random Forests.
+Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples.
+A system is deemed to be effective if the model can discriminate between such samples with low error.
 (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].)
-We observe that the vast majority of proposals do not directly measure entropy, which is a crucial factor in determining whether sensor data cannot be easily predicted by an adversary.
-Some work has attempted to assess the entropy of sensor signals, but this represents a minority.
+Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a minority of work in the literature.
 T2Pair by Li et al.
-[23] yielded 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al.
-[24] reporting 51–54 bits of Shannon entropy.
+[21], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy.
+A refinement by Wu et al.
+[22] reported 51–54 bits of Shannon entropy.
 Even in the best cases, this is low relative to modern cryptographic standards.
 [^2] See AIS 20/31 [23]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator.
-Lastly, we point to adjacent work that has questioned the utility of mobile sensors.
+We also point to adjacent work that has questioned the utility of mobile sensors.
 Markantonakis et al.
@@
 [5] and Shepherd et al.
-[6], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints for NFC-based contactless transactions.
-Using a 500ms window for data acquisition, as stipulated by the EMV payment protocol, 0.179–0.246 EER was reported depending on the given sensor combination.
+[6], presented a reproducibility study of mobile sensors when deployed under a 500ms time constraint for NFC-based transactions as specified by the EMV payment protocol.
+Here, 0.179–0.246 EER was reported depending on the given sensor combination.
 Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
@@
 In earlier work, Voris et al.
-[22] investigated the use of accelerometers as true random number generators (TRNGs), specifically aboard a WISP RFID tag and a Nokia N97 mobile phone.
-The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, and that stationary movement has the lowest min-entropy.
-Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC) are considered other significant influences on the generated entropy.
+[24] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone.
+The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy.
+Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC), are considered significant influences on entropy generation.
 Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer.
 Lv et al.
-[26] analyse three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer.
-Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method used.
+[25] analysed three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer.
+Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method.
 Krhovják et al.
-[7] examine the entropy of image and audio data collected from mobile phone cameras and microphones respectively.
-Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropy values of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera).
+[7] examined the entropy of image and audio data collected from mobile phone cameras and microphones respectively.
+Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera).
 Hennebert et al.
-[27] present an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
-The work uses a single-sensor analysis, presenting min-entropy values between 0–7.85.
-Motion sensors, e.g.
-accelerometer and vibration sensors, yielded the highest entropy.
-Mobile sensors have also been proposed as entropy sources in low-cost RNG designs.
+[26] presented an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
+A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g.
+accelerometer and vibration sensors, produced the highest entropy.
+Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices.
 Suciu et al.
-[28] propose using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors.
-Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [29] suite, but no precise entropy values are presented.
+[27] proposed using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors.
+Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values were presented.
 Wallace et al.
-[30] explore the use of accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources.
-Mobile sensors are inherently opportunistic entropy sources, contrasting with dedicated hardware like ring oscillators, those build on Johnson-Nyquist thermal noise, or quantum-based generators (e.g.
+[29] explored an RNG design using the accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources.
+Results are presented from a non-standard entropy evaluation using 37 Android devices.
+Sensors within existing work serve as inherently opportunistic entropy sources, i.e.
+in contrast with dedicated TRNGs using ring oscillators, Johnson-Nyquist thermal noise, and quantum phenomena (e.g.
 see [30]).
-Mobile sensors depend heavily on user behaviour and environmental, introducing biases and correlations absent in controlled entropy sources.
-Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic or model evaluation metrics [8, 11].
-Such approaches do not account well for skewed distributions and other biases in the underlying data.
-This contrasts with typical entropy measures used in RNG analyses or in the study of authentication systems [32].
+Mobile sensors depend heavily on user behaviour and environmental, which could result in biases and correlations that are absent in controlled entropy sources.
+Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics [8, 11].
+Such approaches do not account well for skewed distributions and other biases in the underlying data that affect predictability.
+Contrast this with typical measures used in the area of randomness testing and authentication [31, 32, 30, 33, 23].
 For instance, NIST SP800-90B [33] and AIS 20/31 [23] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source.
@@
 5.1 Comparison with Related Work
-Prior studies have claimed that mobile sensors can serve as a robust source of security or randomness.
-Much work has relied on machine-learning-based metrics as a proxy to evaluate such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics.
-Li et al.[23] report 32–38 bits of Shannon entropy for some motion-based secure pairing schemes, and Wu et al.[24] estimate 51–54 bits.
-However, dedicated entropy analyses frequently demonstrate that individual sensors confer very little min-entropy, especially under worst-case assumptions [26, 27, 22, 7].
-Our results, derived from a systematic analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
+Prior studies have claimed or implied that mobile sensors are suitable data sources for security-critical applications.
+Much work has relied on model evaluation metrics as a proxy for evaluating such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics [22, 21].
+However, analyses have shown that individual sensors confer very little entropy, especially under worst-case assumptions [25, 26, 24, 7].
+Our results, derived from an analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
 Single sensors yield low entropy.

### v6

@@
-Newcastle-upon-Tyne, UK
+Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security
 Abstract
 Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication.
-However, the foundational assumption that these signals provide sufficient entropy remains under-explored.
+However, the foundational premise that these signals provide sufficient entropy remains under-explored.
 In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple application contexts.
-Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher.
+Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings.
 We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy.
 This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities.
-These results call into question the widely held assumption that adding more sensors inherently yields higher security.
-We ultimately caution against relying on raw sensor data as a primary source of randomness.
+These results raise the serious risk of attacks that exhaustively search the space of possible sensor measurements.
+Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
 keywords:
-Entropy Sensors Mobile security
+Entropy , Sensors , Mobile security
 1 Introduction
@@
 And, secondly, to what extent do multi-modal sensor combinations provide security gains?
-Our analysis reveals systemic limitations.
-Commodity sensors exhibit significant biases, with gyroscopes yielding min-entropy as low as 0.96 bits.
-While multi-modal sensor usage somewhat addresses these issues, non-uniform distributions and inter-sensor correlations reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy.
-These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using raw sensor data as a dependable entropy source.
+Understanding the underlying entropy is important: even if different sensors are combined, fused or otherwise transformed, it does not fundamentally improve the quantity of entropy, or unpredictability, inherent in such signals.
+This paper investigates those concerns.
+Our analysis reveals systemic limitations: commodity sensors exhibit significant biases of between 3.408–4.483 bits of min-entropy (5.584–9.266 bits of Shannon entropy on average).
+In this paper, we analyse 25 different sensors compared to a far smaller number explored in related work, i.e.
+[15] (1 sensor), [16] (10), [7] (2), and [17] (3).
+Furthermore, to the best of our knowledge, we also present the first multi-modal entropy analysis at this scale.
+We find that, while multi-modal sensor usage confers some benefits, non-uniform distributions and inter-sensor correlations significantly reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy.
+These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using sensor data as a dependable entropy source.
 Our contributions are as follows:
-  • We introduce the first systematic approach to evaluating sensor entropy across diverse modalities and datasets using various entropy metrics (max, Shannon, collision, and min-entropy).
+  • 1.
+    We introduce the first systematic approach to evaluating sensor entropy across such a comprehensive range of modalities and datasets using various entropy metrics (max, Shannon, collision, and min-entropy).
+  • 2.
     We empirically demonstrate how inter-sensor correlations and biases erode entropy, casting doubt on the proposition that using multiple sensors adds substantially to security.
+  • 3.
+    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the (joint) measurement space.
+This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
+  • 4.
     Ultimately, we advise against relying on commodity sensors as sources of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
@@
 2 Background
-This section discusses sensor-based applications primitives, critiques existing approaches, and formalises the entropy metrics underpinning our analysis.
+This section discusses sensor-based applications, critiques existing approaches, and formalises the entropy metrics underpinning our analysis.
 2.1 Mobile Sensor-based Security Applications
@@
 T2Pair by Li et al.
-[21], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy.
-A refinement by Wu et al.
-[22] reported 51–54 bits of Shannon entropy.
+[24], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy; a refinement by Wu et al.
+[25] reported 51–54 bits.
 Even in the best cases, this is low relative to modern cryptographic standards.
 [^3] See AIS 20/31 [26]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator.
-We also point to adjacent work that has questioned the utility of mobile sensors.
+We also point to work that has questioned the utility of mobile sensors in time-critical domains.
 Markantonakis et al.
@@
 Hennebert et al.
-[26] presented an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
+[16] presented an analysis of 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
 A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g.
@@
 We use the following definitions and notation throughout this work.
-Definition 1 (Rényi Entropy)
+Definition 2.1 (Rényi Entropy).
 Let $X$ be a discrete random variable taking values in a set $\mathcal{X}$ with probability mass function $p(x)$. The Rényi entropy of order $\alpha$ ($\alpha>0$, $\alpha\neq 1$) is defined as
@@
 We also rely on joint entropy for assessing the entropy of multiple random variables.
-Definition 2 (Joint Rényi Entropy)
+Definition 2.2 (Joint Rényi Entropy).
 Let $X_{1},X_{2},\dots,X_{n}$ be discrete random variables that jointly take values in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$, with joint probability mass function $p(x_{1},x_{2},\dots,x_{n})$.
@@
 To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we also define the Kullback-Leibler divergence.
-Definition 3 (Kullback-Leibler (KL) Divergence)
+Definition 2.3 (Kullback-Leibler (KL) Divergence).
 Given a true probability distribution $P(x)$ of a random variable and an approximate or reference distribution $Q(x)$, the KL divergence is defined as follows:
@@
 3.1 Threat Model and Assumptions
-We consider an adversary whose primary objective is to compromise sensor-based key generation or authentication schemes by exploiting the inherent biases or predictable patterns within sensor data.
-We assume the adversary has full or near-full knowledge of the probability distribution of sensor outputs.
-They may obtain this through statistical modeling of publicly available datasets, prolonged observation of target users, or eavesdropping on sensor data streams.
-We exclude capabilities such as fault-injection and other hardware attacks, e.g.
-see [34].
-While such attacks could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard this out-of-scope for this work.
-Furthermore, rather than brute-forcing all possible sensor outputs, we assume a more strategic approach.
-The adversary uses its knowledge of the sensor’s distribution to prioritise the most probable outcomes.
-By iterating guesses in order of descending probability, they minimise their expected effort.
+We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication).
+The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets.
+We assume the attacker focuses on predicting or guessing the sensor outputs by prioritising the most probable values first, exploiting any biases in the distribution of sensor measurements.
+As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough.
+We exclude capabilities such as fault injection and other hardware attacks (see [34]).
+While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as out-of-scope in this work.
 This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically.
-Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist knowledge-driven attacks.
+Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
 3.2 High-level Methodology
@@
 Ultimately, we narrowed our scope to four datasets that offer diverse usage contexts, consistent sampling rates, and documented sensor modalities:
+  • 1.
     UCI-HAR [39]: A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities.
 Data includes triaxial accelerometer and gyroscope signals.
+  • 2.
     University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz from an Huawei Mate 9 smartphone.
 The publicly available SHL Preview dataset is used, comprising three recording-days per user (59 hours of data in total).
-To scope this study, we use the dataset from the handheld mobile phone as a good fit with related work (see §2).
+To scope this study, we use the dataset from the handheld mobile phone as a good fit with related work.
+  • 3.
     Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés).
 The dataset encompasses accelerometer, gyroscope, and environmental readings taken in realistic payment scenarios.
+  • 4.
     PerilZIS [42]: Collected at 10 Hz from a Texas Instruments SensorTag, a Samsung Galaxy S6, and a Samsung Galaxy Gear, this dataset spans multiple zero-interaction security use cases in an office environment.
-These four datasets provide a variety of sensor types, user activities, and sampling rates, allowing us to explore how intrinsic biases and correlations manifest across different real-world scenarios.
+These four datasets provide a variety of sensor types, user activities, and sampling rates, allowing us to explore how intrinsic biases and correlations manifest across different scenarios.
 Next, we detail how we preprocess and aggregate this data to form global distributions for our entropy analyses.
@@
 4.2 Single Sensors
-Given the biases discussed above, it is inevitable that some sensor readings will exhibit relatively high predictability.
-To quantify this, we calculate individual-sensor entropies across multiple datasets.
-The results are given in Table 1.
-For multi-dimensional modalities (e.g.
-triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al.
-[24].
-We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors.
-Rather, the authors have already preprocessed triaxial data into its vector magnitudes, i.e.
-$\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$.
-We give this as “X.Mag” for a given sensor X.
-For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this new synthetic modality.
-Several clear patterns emerge from Table 1.
-Some sensors (e.g.
-certain accelerometer axes in SHL or PerilZIS) exhibit moderate min-entropies (4–6 bits).
-Other sensors, particularly gyroscope axes, (see UCI-HAR), show values below 3 bits, indicating high predictability in their most frequent readings.
-Shannon entropy values ($H_{1}$) can be fairly high (up to 10 bits in some cases), whereas min-entropy ($H_{\infty}$) is often much lower.
-This gap reflects distributions where a few outcomes dominate, thereby driving worst-case unpredictability down even if the average-case picture is more favorable.
-Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own.
-In the next section, we examine whether combining multiple modalities can meaningfully increase this worst-case unpredictability or whether correlated biases persist across different sensor streams.
+Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data for that dataset and modality.
 Dataset
@@
 S.D. 0.207 0.904 1.461 1.574 1.993 1.148 1.115 1.018 1.612 1.227 1.474 1.379 2.312 0.900 1.526 1.266
-  • *: The authors preprocess triaxial data into its vector magnitude.
-Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data.
+Given the biases discussed above, it is inevitable that some sensor readings will exhibit relatively high predictability.
+To quantify this, we calculate individual-sensor entropies across multiple datasets.
+The results are given in Table 1.
+For multi-dimensional modalities (e.g.
+triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al.
+[15].
+We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors.
+Rather, the authors have already preprocessed triaxial data into its vector magnitudes, i.e.
+$\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$.
+We give this as “X.Mag” for a given sensor X.
+For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this new synthetic modality.
+Several clear patterns emerge from Table 1.
+Some sensors, such as certain accelerometer axes in SHL or PerilZIS, exhibit moderate min-entropies of 4–6 bits.
+Other sensors, particularly gyroscope axes (see UCI-HAR) show values below 3 bits, indicating high predictability in their most frequent readings.
+Shannon entropy values ($H_{1}$) can be fairly high (up to 10 bits in some cases), whereas min-entropy ($H_{\infty}$) is often much lower.
+This gap reflects distributions where a few outcomes dominate, thereby driving worst-case unpredictability down even if the average-case picture is more favorable.
+Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own.
+In the next section, we examine whether combining multiple modalities can meaningfully increase this worst-case unpredictability or whether correlated biases persist across different sensor streams.
 4.3 Multi-modal Sensors
@@
 This section will examine that claim.
-A naïve approach to analysing the entropy of multiple sensors is to first assume that each sensor is statistically independent, yielding the simplified relation $H(X_{1},X_{2},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
-However, real-world sensor outputs often exhibit significant interdependence.
-For example, the linear acceleration, rotation vector, and gravity sensors in the Android ecosystem can be implemented in software using data from accelerometers and gyroscopes, creating a source of correlation [44].
-Figure 2 shows the correlation matrices of the modalities across our four datasets, revealing that many modalities are far from independent.
 (a) UCI-HAR
@@
 Figure 2: Sensor correlation matrices for each dataset.
-4.3.1 Complexity challenges.
-Unfortunately, directly estimating the joint entropy of discrete sensor data becomes computationally infeasible as the number of modalities increases due to the curse of dimensionality.
-Specifically, for $n$ modalities, where each modality $X_{i}$ is discretised into $b_{i}$ bins (with $\{b_{1},b_{2},\dots,b_{n}\}\subseteq B$), the total number of joint bins scales as $\prod^{n}_{i=1}b_{i}$.
+A naïve approach might add Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
+However, this requires that $X_{i}$ are statistically independent.
+In reality, mobile sensors often exhibit strong dependencies.
+For instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices [44].
+As a result, these modalities cannot be treated as independent random variables.
+Figure 2 illustrates how multiple sensors in each dataset correlate: some pairs are nearly perfectly aligned (correlation close to ±$\pm 1$), which drastically reduces their combined unpredictability.
+High correlations invalidate the simplistic additive model of entropy.
+Even if multiple modalities individually appear to have moderate unpredictability, overlapping probability distributions may limit the overall joint entropy.
+In the next subsections, we discuss why straightforward joint-entropy calculations are computationally intractable at scale, before describing how Chow–Liu trees enable a practical approximation of higher-dimensional entropy.
+4.3.1 Complexity Challenges
+Computing the exact joint probability distribution and joint entropy of multiple sensors can quickly become prohibitively expensive.
+Let each of the $n$ sensor modalities be discretised into $b_{i}$ bins.
+Then, the joint distribution has $\prod_{i=1}^{n}b_{i}$ distinct states, an enormous state space once $n$ and $b_{i}$ grow.
 Applying Freedman–Diaconis binning rules typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially.
-Consequently, computing a full joint entropy estimation, as in Eq.
-6, becomes intractable.
-Another concern is the combinatorial explosion of sensor combinations.
-Let us assume that one has 19 modalities, as in Table 1, then the total number of combinations to analyse is equal to $202^{19}-$ 20, skipping single sensors and the empty set.
-For even moderately large datasets, iterating over every sensor subset and computing its full joint entropy is prohibitively expensive.
-Addressing these problems is challenging.
-One mitigation is to reduce the number of bins per modality.
-However, this inevitably risks oversimplifying the distribution and underestimating its joint entropy.
+Moreover, even before enumerating states, selecting which sensors to combine can itself involve $2^{n}-(n+1)$ subsets, skipping single-sensor subsets and the empty set.
 Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset.
-Further scaling became infeasible.
-We also observed that imposing these smaller bin counts (rather than the thousands implied by Freedman–Diaconis) reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1.
-4.3.2 Chow-Liu approximation.
-To address the problem of high-dimensional entropy estimation, we adopt the Chow-Liu algorithm [45] to approximate the joint probability distribution of multiple sensors using a tree-structured Bayesian network.
-Specifically, Chow-Liu constructs a maximum-weight spanning tree $\pi$ over the $n$ modalities, where the edge weights correspond to the mutual information between pairs of variables.
-This tree structure minimises the Kullback-Leibler divergence (Def.
-2.3) between the true joint distribution and the tree-structured approximation.
-The joint probability is then estimated as:
+Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates.
+Further experiments confirmed that limiting the bin numbers reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1.
+We therefore sought an alternative strategy that balances accuracy with tractable computation.
+4.3.2 Chow-Liu Approximation
+To handle these scaling issues, we adopt Chow–Liu trees [45], which approximate high-dimensional joint distributions using a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
+Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies.
+This approach minimises the Kullback–Leibler divergence (Def.
+2.3) between the true multivariate distribution and the resulting tree-based approximation as follows:
 $p_{\pi}(x_{1},\dots,x_{n})=p(x_{r})\prod_{i\neq r}p(x_{i}\;|\;x_{\pi(x)})$ (9)
 Where $\pi(i)$ denotes the parent of $X_{i}$ in the tree, and $r$ is the tree’s root node.
-This model efficiently captures the dominant pairwise dependencies, which is particularly beneficial when exploring the combinatorial relationships of different modalities.
-The use of Chow-Liu trees was also suggested by Buller and Kaufer [35] for estimating the entropy of multivariate data sources where the range of possible values is high.
+Chow–Liu trees are acyclic, singly connected structures: each node has at most one parent where one can traverse the tree to accumulate probabilities between pairwise dependencies.
+This significantly reduces computation time compared to naïve enumeration of the full joint measurement space.
+The use of Chow-Liu trees was proposed by Buller and Kaufer [35] for estimating the entropy of multivariate data sources where the range of possible values is high.
 In our Python implementation, we use the pgmpy [46] library’s TreeSearch module.
-With the inferred Bayesian network in hand, we can compute the max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$) and min-entropies ($H_{\infty}$) without resorting to summation over the full exponential state space.
-Chow–Liu’s acyclic, singly connected structure is crucial: each node has at most one parent, so one can traverse the tree top-down (from root) or bottom-up to accumulate either sums of probabilities (e.g.
-for collision entropy) or maximum probabilities (e.g.
-min-entropy).
-This significantly reduces computation time compared to naïve enumeration.
-Our entropy analysis framework systematically evaluates the joint entropy over all valid combinations of sensor modalities.
+Practically, for each sensor subset, we:
+  1. 1.
+    Discretise each sensor’s readings via Freedman–Diaconis binning.
+  2. 2.
+    Build a Chow–Liu tree from the mutual information of each sensor pair, selecting edges to form a spanning tree.
+  3. 3.
+    Traverse the resulting tree to estimate max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$), and min-entropy ($H_{\infty}$) without enumerating the full exponential state space.
+Our framework evaluates the joint entropy over all sensor combinations.
 The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module.
-Each subset is independently analysed: quantising the data, inferring a Chow–Liu tree, and computing the suite of entropy metrics.
 Processing all four datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
-4.3.3 Results.
-The results are given in Tables 2 to 5, listing the top 10 sensor combinations for each dataset in descending order of min-entropy.
-The full sensor set consistently achieves the highest entropy values.
-This behavior is anticipated: adding more modalities increases the theoretical upper bound on entropy (e.g., $H_{0}$), and it does not lower the min-entropy.
-We observe that removing certain modalities often has a negligible impact on worst-case unpredictability.
-For example, in the SHL dataset (Table 5), while the full sensor suite attains $H_{0}=158.601$ bits and $H_{1}=82.301$ bits, the min-entropy remains at a constant $H_{\infty}=21.289$ bits across both the full sensor set and the top-performing subsets.
-This consistency indicates that, although additional sensors expand the total support (reflected in higher $H_{0}$), they do not necessarily enhance worst-case unpredictability ($H_{\infty}$).
-Similarly, in Table 3 for the Relay dataset, the combination (Acc., Gyro., Light, Lin.
+4.3.3 Results
+Tables 25 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset.
+As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy.
+However, min-entropy ($H_{\infty}$) remains stubbornly low.
+For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ (Shannon) but saturates at only 21 bits of $H_{\infty}$.
+Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all.
+For the Relay dataset in Table 3, the combination (Acc., Gyro., Light, Lin.
 Acc., Mag., Rot.
 Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits.
-Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$.These observations suggest that carefully chosen sensor subsets can yield nearly optimal min-entropy while reducing the number of modalities and associated overhead (e.g.
-battery consumption).
+Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$.
+This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
+The results imply that many sensors contribute redundant information, showing a fundamental limitation of multi-modal data in real-world devices.
+Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof.
+We note that standards use min-entropy as a safer, worst-case metric nowadays [26, 33].
+It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first.
+To the best of our knowledge, the entropy collapse brought about by highly correlated sensor modalities has not been before in existing work.
 We provide the full sensor combination results in our open-source repository.
@@
 We then discuss potential mitigations such as randomness extractors, outline inherent limitations of our study and sensor-based approaches, and propose avenues for future enhancements.
-5.1 Comparison with Related Work
+5.1 Discussion
 Prior studies have claimed or implied that mobile sensors are suitable data sources for security-critical applications.
 Much work has relied on model evaluation metrics as a proxy for evaluating such claims [13, 14, 4, 2], with a smaller subset using more established entropy metrics [25, 24].
-However, analyses have shown that individual sensors confer very little entropy, especially under worst-case assumptions [25, 26, 24, 7].
+However, analyses have demonstrated that individual sensors confer very little entropy, especially under worst-case assumptions [17, 16, 15, 7].
 Our results, derived from an analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
 Single sensors yield low entropy.
-In line with prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits).
+Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits).
 Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g.
 some single-sensor readings yield only 1–3 bits of min-entropy).
-This highlights a major gap between average-case unpredictability and worst-case security, which is an important distinction that many earlier evaluations do not fully capture.
+This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not fully capture.
 Multiple sensors improve worst-case entropy, but not dramatically.
@@
 In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution.
-Min-entropy is critical in scenarios where an adversary can guess the most likely states first.
-Overall, our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches.
-We estimate that even the most complicated multi-modal combination provide relatively little entropy.
+Our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches.
+We estimate that even the most complicated multi-modal combination provide relatively little worst-case entropy.
 It is important to note that our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further.
@@
 They might be useful for tasks where strong assurances are not required, e.g.
-simple proximity checks; however, the suitability of sensor data as a robust source of unpredictability remains highly doubtful.
-5.1.1 Can Extractors Salvage Weak Sources?
+simple proximity checks; however, relying on sensor data as robust entropy sources is fraught with security risks.
+What can be done?
 One might hope that cryptographic extractors (e.g.
-Von Neumann extractors or more advanced schemes [47, 48, 49]) could mitigate low-entropy sensor data.
-Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of randomness beyond the source’s intrinsic min-entropy.
-That is, if the combined sensor data only provides around 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more.
-From a security standpoint, this means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance).
-Thus, while extraction can help if the original data has enough min-entropy to meet system requirements, it cannot elevate a weak source—one that fundamentally lacks sufficient min-entropy—to the higher security thresholds often demanded by rigorous standards.
+Von Neumann extractors or more advanced schemes [47, 48, 49]) could make low-entropy sensors suitable for usage.
+Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of unpredictability beyond the source’s intrinsic min-entropy.
+If combined sensor data provides, say, 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more.
+This means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance).
+Extraction enhances the quality of the original source if it has enough min-entropy, but it cannot elevate a source that fundamentally lacks it.
+Another avenue is to introduce additional unpredictability through user interaction or deliberate environmental perturbation.
+For example, requiring the user to perform a random shaking gesture during a pairing protocol could inject extra entropy into the sensor readings.
+Prior studies have found that deliberate device movements can, indeed, increase the usable entropy of motion sensors, albeit only by a modest amount (on the order of a few bits) [15, 17].
+While this boost is non-negligible, it remains far below the dozens or hundreds of bits typically desired for security-critical applications.
+One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudorandom number generator (CSPRNG).
+By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al.
+[27] and Wallace et al.
+[29]).
+However, such an approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings, such as matching motion traces for device pairing or zero-interaction authentication [42, 13, 2, 3, 4, 1].
+Here, the desired property is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices.
+We have shown that sensor signals inherently have low entropy, even when using multiple sensors simultaneously, exposing a fundamental security flaw in such designs.
+We see no straightforward remedy for the entropy shortfall of mobile sensors.
+Extractors, user-assisted randomness, and entropy pools address different problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution.
+Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities.
+Sensor data and its derivatives may still be appropriate for low-security purposes.
+However, relying on it alone for security applications is highly inadvisable due to the serious risk of adversaries who exploit statistical biases to reduce the effective search space.
 5.2 Limitations
-Despite our comprehensive analysis, this work has some limitations.
+Despite our detailed analysis, this work has some notable limitations.
 Firstly, while we analyse four large datasets, they do not fully capture contexts all possible contexts.
@@
 The effectiveness of countermeasures remains an open research challenge.
-Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, and ultimately pave the way for safer and more robust designs in mobile-device security.
+Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs in mobile device security.
 In future work, we consider that a dynamic analysis is important to assess the stationarity issues with sensor data, where entropy varies according to user behavior or environment.
@@
 Overall, while sensor-based data can reach limited levels of entropy under favourable conditions, the road to making such sources systematically secure and robust is long.
-The key takeaway is that researchers and engineers should resist the temptation to treat sensor readings as bona fide random bits.
-Substantial work is needed before sensors can be considered as appropriate sources of randomness for security-critical applications.
-6.0.1 Acknowledgments
+The key takeaway is that substantial work is needed before sensors can be considered appropriate for security-critical applications.
+Acknowledgments
 Carlton Shepherd has received funding from the UK EPSRC ‘Chameleon’ project (EP/Y030168/1). The authors would like to thank Darren Hurley-Smith for insightful conversations on the topic.

### v7

@@
 However, the foundational premise that these signals provide sufficient entropy remains under-explored.
-In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple application contexts.
+In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple contexts.
 Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings.
@@
 This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities.
-These results raise the serious risk of attacks that exhaustively search the space of possible sensor measurements.
+Our results demonstrate that adversaries may feasibly predict sensor signals through an exhaustive, brute-force exploration of the entire measurement space.
 Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
@@
 1 Introduction
-Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity.
-This rich sensor data has enabled applications from activity recognition to context-aware computing.
+Modern mobile devices are equipped with various embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity.
+This sensor data has enabled myriad applications from activity recognition to context-aware computing.
 More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication.
@@
   • 3.
-    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the (joint) measurement space.
+    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the space of possible measurements.
 This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
   • 4.
-    Ultimately, we advise against relying on commodity sensors as sources of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
+    Ultimately, we do not recommend relying on commodity sensors as a source of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
 The rest of this paper is organised in the following way: §2 discusses sensor-based security mechanisms and established entropy metrics.
@@
 We exclude capabilities such as fault injection and other hardware attacks (see [34]).
-While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as out-of-scope in this work.
+While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work.
 This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically.
@@
     Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés).
-The dataset encompasses accelerometer, gyroscope, and environmental readings taken in realistic payment scenarios.
+The dataset contains accelerometer, gyroscope, and environmental readings from devices involved in emulated payment scenarios.
   • 4.
@@
 The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module.
-Processing all four datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
+Processing all datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
 4.3.3 Results
@@
 This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
-The results imply that many sensors contribute redundant information, showing a fundamental limitation of multi-modal data in real-world devices.
+(a) UCI-HAR
+(b) PerilZIS
+(c) Relay
+(d) SHL
+Figure 3: Entropy values versus the number of joint sensor modalities.
+[…]
+The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality.
 Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof.
-We note that standards use min-entropy as a safer, worst-case metric nowadays [26, 33].
+We note that standards implore the use of min-entropy as a safer, worst-case metric [26, 33].
 It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first.
-To the best of our knowledge, the entropy collapse brought about by highly correlated sensor modalities has not been before in existing work.
-We provide the full sensor combination results in our open-source repository.
+To the best of our knowledge, the entropy collapse from highly correlated sensor modalities has not been elicited in existing work.
+We provide the full sensor combination results in our open-source repository.2
 Table 2: Top 10 best-performing sensor combinations (UCI-HAR; in bits).
@@
 While this boost is non-negligible, it remains far below the dozens or hundreds of bits typically desired for security-critical applications.
-One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudorandom number generator (CSPRNG).
+One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudo-random number generator (CSPRNG).
 By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al.
@@
 [29]).
-However, such an approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings, such as matching motion traces for device pairing or zero-interaction authentication [42, 13, 2, 3, 4, 1].
-Here, the desired property is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices.
-We have shown that sensor signals inherently have low entropy, even when using multiple sensors simultaneously, exposing a fundamental security flaw in such designs.
+However, this approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings.
+For example, in matching similar motion traces or environmental conditions of two neighbouring devices [42, 13, 2, 3, 4, 1].
+The desired property here is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices.
+Using CSPRNG-derived signals would make it difficult to distinguish whether signals originated from co-located devices using similarity-based metrics.
+This exposes a fundamental security paradox in existing work.
+We have shown that sensors have insufficient entropy, even when using multiple ones simultaneously.
+Yet, methods to increase that entropy would compromise the usefulness of schemes that rely on the similarity of one or more sensor signals.
 We see no straightforward remedy for the entropy shortfall of mobile sensors.
-Extractors, user-assisted randomness, and entropy pools address different problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution.
+Extractors, user-assisted randomness, and entropy pools address diHoweverfferent problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution.
 Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities.

### v8

@@
 However, the foundational premise that these signals provide sufficient entropy remains under-explored.
-In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple contexts.
-Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings.
-We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy.
-This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities.
-Our results demonstrate that adversaries may feasibly predict sensor signals through an exhaustive, brute-force exploration of the entire measurement space.
-Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
+In this work, we systematically analyse the entropy of mobile sensor data using four datasets from multiple application contexts (UCI-HAR, SHL, Relay, and PerilZIS).
+Using direct computation and estimation, we report entropy values—max, Shannon, collision, and min-entropy—for an exhaustive range of sensor combinations.
+We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, such as for authentication and key generation systems, even when many sensors are combined.
+In particular, we observe an alarming divergence between average-case Shannon entropy and worst-case min-entropy.
+Single-sensor mean min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher.
+We further show that correlations between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy.
+This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits when combining 20+ sensor modalities.
+Our results reveal that adversaries may feasibly predict sensor signals through an exhaustive exploration of the measurement space.
+Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security.
+Ultimately, we strongly urge caution when relying on mobile sensor data for security applications.
 keywords:
@@
 1 Introduction
-Modern mobile devices are equipped with various embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity.
-This sensor data has enabled myriad applications from activity recognition to context-aware computing.
-More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication.
-A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference.
-Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11].
-Despite these advancements, fundamental issues remain: multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated.
-Many existing studies rely on heuristic assessments or machine learning classifiers, e.g.
-[4, 13, 14, 2, 12], that do not address critical security questions.
-That is, firstly, how much entropy do sensors truly provide?
-And, secondly, to what extent do multi-modal sensor combinations provide security gains?
-Understanding the underlying entropy is important: even if different sensors are combined, fused or otherwise transformed, it does not fundamentally improve the quantity of entropy, or unpredictability, inherent in such signals.
-This paper investigates those concerns.
-Our analysis reveals systemic limitations: commodity sensors exhibit significant biases of between 3.408–4.483 bits of min-entropy (5.584–9.266 bits of Shannon entropy on average).
-In this paper, we analyse 25 different sensors compared to a far smaller number explored in related work, i.e.
-[15] (1 sensor), [16] (10), [7] (2), and [17] (3).
-Furthermore, to the best of our knowledge, we also present the first multi-modal entropy analysis at this scale.
-We find that, while multi-modal sensor usage confers some benefits, non-uniform distributions and inter-sensor correlations significantly reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy.
-These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using sensor data as a dependable entropy source.
+Modern mobile devices expose rich motion and ambient signals from accelerometers, gyroscopes, magnetometers, and more.
+Beyond activity recognition and context inference, e.g.
+[1, 2, 3, 4], a long line of work proposes using these signals for security-critical tasks.
+This includes cryptographic key generation, zero-interaction device pairing, and continuous authentication [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
+Traditional key generation protocols [8, 18] have historically relied on shared motion patterns in order to establish secure communication between co-located devices.
+Meanwhile, other proposals have incorporated environmental phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
+Despite these advances, a crucial assumption remains under-examined: that sensor data provides sufficient entropy to thwart adversarial inference.
+Many existing studies rely on heuristic assessments or model classsification metrics, e.g.
+[21, 7, 6, 9, 22, 19, 20], that do not address two basic questions: (1) how much entropy do sensors actually provide, and (2) to what extent do multimodal sensor combinations improve entropy?
+For the latter, multimodal sensing has been advocated to counter sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated.
+The security implications of low-entropy sensor data can be direct and severe.
+If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it.
+For such reasons, modern security standards provide important quantitative thresholds for what can be considered ‘secure’.
+For instance, NIST SP 800-57 [23] states that security levels below 80 bits are broken for long-term use, with 112 bits as the current minimum and 128 bits recommended for security keys.
+Standards that govern dedicated entropy sources are even stricter: NIST SP 800-90B [24] expects entropy sources used for cryptographic random number generators (RNGs) to approach full per-bit entropy, [^2] More specifically, $\approx$1 bit ($\epsilon\leq 2^{-32}$ per-bit) of min-entropy per output bit.
+while AIS 20/31 [25] requires an RNG’s internal state to possess over 240 and 250 bits of min-entropy and Shannon entropy respectively.
+As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls short of these standards.
+In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—[26] (1 sensor), [27] (10), [17] (2), [28] (3)—across multiple public datasets.
+Our results reveal structural limits.
+Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets.
+Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically 40–75% below Shannon entropy for the same signals.
+In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be systematically enumerable.
 Our contributions are as follows:
@@
   • 3.
-    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the space of possible measurements.
-This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
-  • 4.
-    Ultimately, we do not recommend relying on commodity sensors as a source of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
-The rest of this paper is organised in the following way: §2 discusses sensor-based security mechanisms and established entropy metrics.
+    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate the space of possible measurements.
+This raises fundamental concerns with respect to security applications that rely on signals from single and multiple mobile sensors.
+Consequently, we do not recommend relying wholly on mobile sensors as entropy sources for security-critical applications.
+The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics.
 §3 explains our experiment design for entropy estimation, including the threat model and dataset selection.
@@
 2 Background
-This section discusses sensor-based applications, critiques existing approaches, and formalises the entropy metrics underpinning our analysis.
-2.1 Mobile Sensor-based Security Applications
-One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [18, 13, 14, 4, 2].
-Mehrnezhad et al.[13] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together.
-The authors posit that “physical tapping causes random but correlated vibrations at both devices, which are hard to forge (or reproduce)” (p.1, [13]).
-The work reports an equal error rate (EER) of 17.65% using a machine learning-based approach.
-Gurulian et al.
-[14] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device.
-Shrestha et al.
-[2] explore four environmental modalities, including temperature and humidity, for proximity detection.
-In this work, single sensors yields 0.733–0.881 F1-score, while combining multiple increases the overall performance to 0.913–0.957.
-Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [19, 20, 21, 8, 9, 22, 10, 11], where sensor data is used to authenticate users passively without explicit interaction (with up to 99% accuracy claimed in some work [20]).
-Sensors have also been used to underpin the security of novel device pairing schemes, whether as a primary [23] or second line of authentication [1].
-These schemes generally rely on detecting similar motion patterns between two devices using joint accelerometer and gyroscope sensor measurements in order to provide evidence of co-location.
-The notion of ‘hardness’ is typically inferred through model evaluation metrics.
-A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g.
-cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2].
-Features are then classified using simple threshold-based or supervised classification models, e.g.
-Support Vector Machines (SVM) and Random Forests.
-Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples.
-A system is deemed to be effective if the model can discriminate between such samples with low error.
-(User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [21, 1].)
-Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a minority of work in the literature.
-T2Pair by Li et al.
-[24], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy; a refinement by Wu et al.
-[25] reported 51–54 bits.
-Even in the best cases, this is low relative to modern cryptographic standards.
-[^3] See AIS 20/31 [26]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator.
-We also point to work that has questioned the utility of mobile sensors in time-critical domains.
-Markantonakis et al.
-[4], building on Gurulian et al.
-[5] and Shepherd et al.
-[6], presented a reproducibility study of mobile sensors when deployed under a 500ms time constraint for NFC-based transactions as specified by the EMV payment protocol.
-Here, 0.179–0.246 EER was reported depending on the given sensor combination.
-Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
-2.2 Sensor Entropy Analyses
-In earlier work, Voris et al.
+This section discusses sensor-based applications and existing approaches, and formalises the entropy metrics underpinning our subsequent analyses.
+2.1 Related Work
+Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context [29, 7, 11, 5, 21, 6, 9]; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret [8, 30, 18, 20]; and (3) continuous authentication, which aims to verify users passively during normal interaction [16, 12, 13, 14, 31, 32, 33, 34].
+Representative designs include tap-to-pair schemes using accelerometers [7], induced shared vibrations [6], and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues [9, 21].
+These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence [8, 30].
+Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based.
+Across the aforementioned application areas, effectiveness is usually inferred from discriminative performance.
+A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g.
+cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1.
+This template appears in numerous systems, including [7, 9, 21, 11, 16, 14, 13, 31, 32, 6].
+For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al.
+[19], Alzubaidi and Kalita [35], and Mayrhofer and Sigg [20].
+While discriminative performance metrics are informative for usability and functionality, they do not quantify entropy of the underlying sources, which is the focus of our analysis.
+Adjacent evidence also highlights practical limits under tight timing constraints.
+For NFC transactions, EMV specifies $\approx$500 ms for proximity checks.
+In that setting, Markantonakis et al.
+[21] (building on [5, 11]) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues.
+Taken together, the literature demonstrates functional feasibility but leaves open the core question of how unpredictable the underlying measurements really are, especially under multimodal fusion.
+This motivates our entropy-centric evaluation in the remainder of the paper.
+Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature.
+Specifically, T2Pair by Li et al.
+[15] reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al.
+[36] reports 51–54 bits.
+Even in the best cases, these values are low relative to modern cryptographic standards; moreover, we are aware of no work using standard min-entropy as a worst-case measure of unpredictability [24, 25, 37] (see comparison tables of state-of-the-art schemes in [35, 18, 20]).
+We address this gap by analysing the entropy of a wide range of mobile sensors, both in isolation and in combination.
+Concurrent work has attempted to assess the entropy bestowed by mobile sensors as a separate line of inquiry.
+Voris et al.
 [26] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone.
-The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy.
-Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC), are considered significant influences on entropy generation.
+The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy.
+Intrinsic noise from the sensor’s circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation.
 Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer.
 Lv et al.
-[17] analysed three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer.
+[28] analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone.
 Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method.
 Krhovják et al.
-[7] examined the entropy of image and audio data collected from mobile phone cameras and microphones respectively.
-Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera).
+[17] examined image and audio data collected from mobile phone cameras and microphones respectively on a Nokia N73 and E-Ten X500 and M700 phones.
+The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera).
 Hennebert et al.
-[16] presented an analysis of 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
+[27] analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
 A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g.
@@
 Suciu et al.
-[27] proposed using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors.
-Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values were presented.
+[38] proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests [39], but without precise entropy estimates.
 Wallace et al.
-[29] explored an RNG design using the accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources.
-Results are presented from a non-standard entropy evaluation using 37 Android devices.
-Sensors within existing work serve as inherently opportunistic entropy sources, i.e.
-in contrast with dedicated TRNGs using ring oscillators, Johnson-Nyquist thermal noise, and quantum phenomena (e.g.
-see [30]).
-Mobile sensors depend heavily on user behaviour and environmental, which could result in biases and correlations that are absent in controlled entropy sources.
-Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics [8, 11].
-Such approaches do not account well for skewed distributions and other biases in the underlying data that affect predictability.
-Contrast this with typical measures used in the area of randomness testing and authentication [31, 32, 30, 33, 26].
-For instance, NIST SP800-90B [33] and AIS 20/31 [26] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source.
+[40] explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices.
+Unlike dedicated TRNGs (e.g.
+ring oscillators, Johnson–Nyquist thermal noise, quantum sources [41]), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources.
+Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability.
+In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 [42, 43, 41, 24, 25].
 Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
+Table 1: Comparative summary of prior sensor entropy analyses.
+Reference Sensorsa Metricsb Results (bits) Context / Use-Case
+Voris et al. [26] Acc. $H_{\infty}$ 3.1–11.4 True random number generation
+Hennebert et al. [27] 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
+Lv et al. [28] Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
+Krhovják et al. [17] Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
+Li et al. [15] Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
+Wu et al. [36] Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
+This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinatons, for security systems
+  • a
+    Acc.: Accelerometer; Gyr.: Gyroscope; Mag.: Magnetometer; Cam.: Camera; Mic.: Microphone.
+  • b
+    $H_{0}$: Max; $H_{1}$: Shannon; $H_{2}$: Collision; $H_{\infty}$: Min-entropy.
+In general, while several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts.
+Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges.
 2.2 Definitions
-We use the following definitions and notation throughout this work.
+We use the following definitions and notation throughout this work. Furthermore, unless stated otherwise, all logarithms are base 2 and entropies are reported in bits.
 Definition 2.1 (Rényi Entropy).
@@
 We draw attention to four special cases of $\alpha$ that are widely used in the literature.
-Firstly, the Hartley (Max) Entropy, given in Eq.
-2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
+First, the Hartley (Max) Entropy (Eq.
+2) is the logarithm of the number of possible outcomes with non-zero probability and serves as an upper bound.
 $H_{0}(X)\equiv\lim_{\alpha\to 0}H_{\alpha}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
-Second is the Shannon Entropy (Eq. 3), which corresponds to the classical definition of entropy in information theory, and is the limit of $H_{\alpha}$ as $\alpha\to 1$.
+Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ as $\alpha\to 1$.
 $H_{1}(X)\equiv\lim_{\alpha\to 1}H_{\alpha}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
-Another case is the Collision Entropy ($\alpha=2$), quantifying the probability of “collisions” of multiple draws from $X$. This is given in Eq. 4.
+Third, the collision entropy ($\alpha=2$) quantifies the probability of drawing the same value twice from $X$:
 $H_{2}(X)\,=\,-\log\Bigl{(}\sum_{x\in\mathcal{X}}p(x)^{2}\Bigr{)}.$ (4)
-$H_{1}$ provides an average-case measure of uncertainty.
+We note that $H_{1}$ provides an average-case measure of uncertainty.
 It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage.
-Min-entropy is thus used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes.
-This is the Min-Entropy (Eq.
-5), i.e.
-the value in the limit $\alpha\to\infty$.
+Min-entropy (Eq.
+5) is used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes; it is the value in the limit $\alpha\to\infty$:
 $H_{\infty}(X)\equiv\lim_{\alpha\to\infty}H_{\alpha}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
-Note that $H_{\alpha}$ is a non-increasing function of $\alpha$, i.e.
-$H_{\infty}(X)\leq H_{2}(X)\leq H_{1}(X)\leq H_{0}(X)$.
+$H_{\alpha}$ is a non-increasing function of $\alpha$, i.e.
+$H_{\infty}(X)\leq\dots\leq H_{1}(X)\leq H_{0}(X)$.
 $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty.
-We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in NIST SP800-90B [33] and AIS 20/31 [26].
+We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards [24, 25].
 We also rely on joint entropy for assessing the entropy of multiple random variables.
@@
 The Rényi entropy of order $\alpha$ ($\alpha>0,\alpha\neq 1$) for these $n$ variables is defined as the following, where the sum is taken over all $(x_{1},\dots,x_{n})$ in $\mathcal{X}_{1}\times\mathcal{X}_{2}\times\cdots\times\mathcal{X}_{n}$:
-$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},\dots,x_{n})}p(x_{1},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
+$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},x_{2},\dots,x_{n})}p(x_{1},x_{2},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
 In the limit $\alpha\to 1$, then $H_{\alpha}(X_{1},X_{2},\dots,X_{n})$ converges to the classical joint Shannon entropy of these $n$ variables.
-To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we also define the Kullback-Leibler divergence.
+To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we define the Kullback-Leibler divergence.
 Definition 2.3 (Kullback-Leibler (KL) Divergence).
@@
 As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough.
-We exclude capabilities such as fault injection and other hardware attacks (see [34]).
+We exclude capabilities such as fault injection and other hardware attacks (see [44]) as well as side-channel attacks such as cross-device correlation attacks [45].
 While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work.
-This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically.
-Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
+This threat model represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically.
+Our goal is to evaluate whether sensor data distributions offer sufficient entropy to resist attacks that search the space of sensor measurement values informed by their statistical properties.
 3.2 High-level Methodology
-We aim to determine the global (i.e., population-level) entropy characteristics of various mobile sensors under ordinary usage conditions, rather than focusing on per-user or scenario-specific differences.
-This choice reflects common real-world deployments, which must accommodate a wide range of behaviors and environments.
+In this work, we focus on estimating population-level (“global”) sensor distributions by pooling all users within a dataset.
+This choice reflects typical deployments in which systems must serve previously unseen users (e.g.
+first-time proximity checks, payment defence mechanisms, or ad-hoc device pairing) and thus cannot assume per-user enrollment or stable user-specific priors.
 Our approach involves five main stages:
   1. 1.
-    We acquire large-scale sensor readings from publicly available datasets that capture diverse user activities and device usage patterns.
-These datasets encompass different motion, environmental, and orientation sensors.
-Detailed descriptions of each dataset are provided at the end of this section.
+    Data acquisition.
+We collect large-scale sensor readings from public datasets spanning diverse activities and device usage.
+Modalities include motion, environmental, and orientation sensors (dataset details appear at the end of this section).
   2. 2.
+    Global distributions.
 We merge sensor readings into a single, global distribution for each sensor modality in each dataset.
@@
   3. 3.
+    Single-sensor entropy analysis.
 From these global distributions, we compute max, Shannon and collusion entropies to measure the best- and average-case uncertainties of sensor outputs, along with the min-entropy to characterise the worst-case unpredictability.
   4. 4.
+    Multimodal entropy estimation.
 Many real-world proposals combine multiple sensor streams to purportedly increase security.
 To assess the impact on worst-case unpredictability, we use Chow-Liu trees to approximate the joint distributions of different sensor modalities.
-This allows us to estimate higher-dimensional entropies without incurring prohibitive computational costs.
-We discuss this in §4.3.
+This allows us to estimate higher-dimensional entropies without incurring prohibitive computational costs (discused in §4.3).
   5. 5.
+    Interpretation.
 Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary.
-We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to similar predictability concerns.
-Throughout this process, we remain mindful of well-documented constraints with NIST SP 800-90B, SP 800-22 [33], and AIS 20/31 [26] in analysing multi-sensor data streams [35, 17].
-Such frameworks were not designed to analyse the joint entropy of complex, multivariate data sources, which is the aim of this work.
-(For example, NIST SP 800-90B focusses on assessing univariate entropy sources with reduced, i.e.
-8-bit, output sizes [35, 33]).
-To begin with, we sought publicly available sensor datasets suitable for analysing motion and environmental data at scale.
-Our search involved broad queries across IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub.
-Several ostensibly “open” datasets either were no longer downloadable or imposed restrictive licensing terms [36, 37, 38].
-Ultimately, we narrowed our scope to four datasets that offer diverse usage contexts, consistent sampling rates, and documented sensor modalities:
+We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to predictability concerns.
+Throughout, we note the constraints of NIST SP 800-90B, SP 800-22 [24], and AIS 20/31 [25] when applied to multi-sensor streams [37, 28].
+These frameworks target univariate sources rather than joint entropy of complex multivariate processes, which is the focus of our work.
+(For example, NIST SP 800-90B focusses on assessing entropy sources with reduced, i.e.
+8-bit, outputs [37, 24]).
+To scope suitable data, we searched IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub.
+Several ostensibly “open” datasets were unavailable for download or had restrictive licensing [46, 47, 48].
+We selected four datasets offering diverse contexts, consistent sampling, and documented modalities:
   • 1.
@@
     University of Sussex–Huawei Locomotion (SHL) [1, 2]: Sampled at 100 Hz from an Huawei Mate 9 smartphone.
-The publicly available SHL Preview dataset is used, comprising three recording-days per user (59 hours of data in total).
-To scope this study, we use the dataset from the handheld mobile phone as a good fit with related work.
+TWe use the publicly available SHL Preview dataset (three recording days per user; 59 hours total) and, for scope alignment with prior work, the handheld-phone subset.
   • 3.
-    Relay [5]: Contains sensor measurements for approximately 1,500 NFC-based contactless transactions, each recorded at 100 Hz across several physical locations (e.g., cafés).
-The dataset contains accelerometer, gyroscope, and environmental readings from devices involved in emulated payment scenarios.
+    Relay [5]: Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g.
+cafés), including accelerometer, gyroscope, and environmental readings in emulated payment scenarios.
   • 4.
-    PerilZIS [42]: Collected at 10 Hz from a Texas Instruments SensorTag, a Samsung Galaxy S6, and a Samsung Galaxy Gear, this dataset spans multiple zero-interaction security use cases in an office environment.
-These four datasets provide a variety of sensor types, user activities, and sampling rates, allowing us to explore how intrinsic biases and correlations manifest across different scenarios.
-Next, we detail how we preprocess and aggregate this data to form global distributions for our entropy analyses.
+    PerilZIS [49]: Collected at 10 Hz from a TI SensorTag, Samsung Galaxy S6, and Samsung Galaxy Gear, spanning multiple zero-interaction security use cases in an office environment.
+These datasets provide varied sensor types, user activities, and sampling rates, enabling us to observe how biases and inter-sensor correlations manifest across scenarios.
+Next, we describe preprocessing and aggregation to form the global distributions used in our entropy analyses.
 4 Entropy Analysis
-In this section, we analyse the intrinsic entropy of sensor data under the threat model described in §3.
-We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our findings for our single- and multi-sensor analyses.
+In this section, we analyse the entropy of sensor data on a single- and multi-modal basis.
+We begin by discussing the challenges in quantising sensor values for discrete-entropy calculations, then present our findings for our single- and multi-sensor analyses.
 4.1 Pre-processing
 A crucial, yet underexplored, issue in prior work (e.g.
-[15, 17, 16, 31]) is how to convert inherently continuous sensor outputs into suitable discrete values for entropy estimation.
-For example, Shannon and min-entropy, as defined in Eqs.
-3 and 5, rely on discrete random variables.
+[26, 28, 27, 42]) is how to appropriately estimate entropy from quantised sensor outputs.
 Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution.
-Yet, a sensor’s advertised resolution (e.g.
-12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform coverage across its range.
-Everyday usage introduces biases and clustering, resulting in some measurements occurring far more frequently than others.
-For instance, UCI-HAR data shows accelerometer readings concentrated in certain areas, and approximately 60% of gyroscope readings hover near zero (see Figure 1).
-Such skew and bias radically diminishes entropy compared to uniformly distributed values.
+Yet, a sensor’s advertised resolution—for instance, 12 bits in the widely used Bosch BMA400 mobile accelerometer [50]—does not imply uniform histograms across its range of values.
+Everyday usage produces highly non-uniform biases and clustering, resulting in some measurements occurring more frequently than others.
+For example, UCI-HAR data shows accelerometer readings concentrated in certain areas, with approximately 60% of gyroscope readings concentrated around zero.
+As an illustration, Figure 1 shows the cumulative probability distributions of some selected modalities across our datasets.
+We note that such skew and bias radically diminishes entropy compared to uniformly distributed values.
 (a) Acc. $x$ axis.
@@
 Figure 1: Global sensor data CDFs – UCI-HAR (a–f) and Relay (g–l) datasets.
-[…]
+Another practical issue is the sparsity of sensor measurements at high resolutions.
+Treating every minute fluctuation e.g.
+$9.001\,\mathrm{m\,s^{-2}}$ vs.
+$9.002\,\mathrm{m\,s^{-2}}$) as distinct outcomes can inflate entropy estimates.
+In real-world applications, it is the ‘similarity’ between measurement signals that is considered useful in existing work; humans cannot reliably reproduce high-precision movements capable of utilising a sensor’s digital resolution at, say, a $0.01\mathrm{ms^{-2}}$ scale.
+We therefore aggregate adjacent codes/values into bins to obtain reliable probability mass functions (pmfs).
+The choice of bin width is non-trivial: classical rules (e.g.
+Sturges, Scott, Doane) encode different assumptions about the underlying distribution, vary in computational cost, and differ in robustness to outliers.
+We adopt the Freedman–Diaconis (FD) rule, a distribution-agnostic, outlier-robust choice that adapts to sample size and spread, [^4] Alternatively, a perceptual/actuation-driven scheme aligned with human reproducibility could be used; we defer this to future work.
+with width $h$ given in Eq.
+8, where $\mathrm{IQR}(x)$ is the interquartile range and $n$ the number of samples.
+Later, in §4.3.4, we show that FD falls within the plateau region as part of a bin sensitivity analysis, yielding stable estimates across datasets while avoiding oversmoothing and sparsity.
 $h=2\cdot\frac{IQR(x)}{n^{1/3}}$ (8)
@@
 UCI-HAR SHL Relay* PerilZIS
-Sensor $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$ $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
+Sensor $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$
 Acc.X 8.488 7.080 5.876 3.729 11.557 8.732 7.487 4.543 13.012 9.292 6.359 3.626
@@
 S.D. 0.207 0.904 1.461 1.574 1.993 1.148 1.115 1.018 1.612 1.227 1.474 1.379 2.312 0.900 1.526 1.266
-Given the biases discussed above, it is inevitable that some sensor readings will exhibit relatively high predictability.
-To quantify this, we calculate individual-sensor entropies across multiple datasets.
+Given the biases discussed above, it is inevitable that some sensor readings will be highly predictable.
+To quantify this, we calculate individual-sensor entropies across datasets.
 The results are given in Table 2.
@@
 [26].
+This approach allows for a fine-grained analysis of directional biases, as movement along one axis may be more constrained or predictable than others in typical usage scenarios.
 We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors.
@@
 $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$.
+This magnitude represents the overall intensity of motion, irrespective of direction.
 We give this as “X.Mag” for a given sensor X.
-For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this new synthetic modality.
+For the sake of completeness and consistency, we compute the magnitude for other datasets, where applicable, and report the entropy values for this synthetic modality.
 Several clear patterns emerge from Table 2.
 Some sensors, such as certain accelerometer axes in SHL or PerilZIS, exhibit moderate min-entropies of 4–6 bits.
-Other sensors, particularly gyroscope axes (see UCI-HAR) show values below 3 bits, indicating high predictability in their most frequent readings.
+Other sensors, particularly gyroscope axes (see UCI-HAR) show values below 3 bits, indicating high predictability.
 Shannon entropy values ($H_{1}$) can be fairly high (up to 10 bits in some cases), whereas min-entropy ($H_{\infty}$) is often much lower.
-This gap reflects distributions where a few outcomes dominate, thereby driving worst-case unpredictability down even if the average-case picture is more favorable.
+This gap reflects distributions where a few outcomes dominate, driving worst-case unpredictability down even if the average-case picture is more favorable.
 Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own.
@@
 4.3 Multi-modal Sensors
-Several sensor-based security proposals [13, 12, 4, 2, 3] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one.
-This section will examine that claim.
+Several sensor-based security proposals [7, 22, 21, 9, 10] assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than one.
+Let us now explore that claim.
+Initially, a naïve approach might add Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
+However, this requires that $X_{i}$ are statistically independent.
+In reality, mobile sensors often exhibit strong dependencies; for instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices [51].
+As a result, these modalities cannot be treated as independent variables.
 (a) UCI-HAR
@@
 (d) SHL
-Figure 2: Sensor correlation matrices for each dataset.
-A naïve approach might add Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
-However, this requires that $X_{i}$ are statistically independent.
-In reality, mobile sensors often exhibit strong dependencies.
-For instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices [44].
-As a result, these modalities cannot be treated as independent random variables.
-Figure 2 illustrates how multiple sensors in each dataset correlate: some pairs are nearly perfectly aligned (correlation close to ±$\pm 1$), which drastically reduces their combined unpredictability.
-High correlations invalidate the simplistic additive model of entropy.
-Even if multiple modalities individually appear to have moderate unpredictability, overlapping probability distributions may limit the overall joint entropy.
-In the next subsections, we discuss why straightforward joint-entropy calculations are computationally intractable at scale, before describing how Chow–Liu trees enable a practical approximation of higher-dimensional entropy.
+Figure 2: Pairwise sensor correlation matrices for each dataset.
+Figure 2 provides a quantitative view of inter-sensor redundancy through Pearson correlation matrices for each dataset.
+Strong correlations, with $|\rho|\approx 1$, indicate high linear dependence.
+Such redundancy is a major contributor to fragility in pairwise combinations.
+Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and provides diminishing returns on min-entropy.
+In other words, while individual modalities may appear moderately unpredictable, the correlations shown in Figure 2 can sharply constrain the overall joint entropy compared to what would be expected if they were independent variables.
 4.3.1 Complexity Challenges
-Computing the exact joint probability distribution and joint entropy of multiple sensors can quickly become prohibitively expensive.
-Let each of the $n$ sensor modalities be discretised into $b_{i}$ bins.
-Then, the joint distribution has $\prod_{i=1}^{n}b_{i}$ distinct states, an enormous state space once $n$ and $b_{i}$ grow.
-Applying Freedman–Diaconis binning rules typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially.
-Moreover, even before enumerating states, selecting which sensors to combine can itself involve $2^{n}-(n+1)$ subsets, skipping single-sensor subsets and the empty set.
+Computing the exact joint probability distribution and joint entropy of multiple sensors quickly becomes infeasible.
+Consider the following: let each of the $n$ sensor modalities be discretised into $b_{i}$ bins, then the joint distribution has $\prod_{i=1}^{n}b_{i}$ distinct states.
+This produces an enormous state space as $n$ and $b_{i}$ grow to the values considered in this paper.
+Applying the FD rule typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially.
+In addition, even before enumerating states, selecting which sensors to combine can itself involve $2^{n}-(n+1)$ subsets, skipping single-sensor subsets and the empty set.
 Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset.
-Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates.
-Further experiments confirmed that limiting the bin numbers reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1.
-We therefore sought an alternative strategy that balances accuracy with tractable computation.
+Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates, as we discussed §4.1.
+In our tests, we found that reducing bin counts reduced single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 2.
+We therefore adopt an alternative strategy that trades some accuracy for tractability.
 4.3.2 Chow-Liu Approximation
-To handle these scaling issues, we adopt Chow–Liu trees [45], which approximate high-dimensional joint distributions using a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
+To handle these scaling issues, we adopt Chow–Liu trees [52] for approximating a high-dimensional joint distribution by a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
 Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies.
 This approach minimises the Kullback–Leibler divergence (Def.
-2.3) between the true multivariate distribution and the resulting tree-based approximation as follows:
+2.2) between the true multivariate distribution and the resulting tree-based approximation as follows:
 $p_{\pi}(x_{1},\dots,x_{n})=p(x_{r})\prod_{i\neq r}p(x_{i}\;|\;x_{\pi(x)})$ (9)
@@
 Chow–Liu trees are acyclic, singly connected structures: each node has at most one parent where one can traverse the tree to accumulate probabilities between pairwise dependencies.
-This significantly reduces computation time compared to naïve enumeration of the full joint measurement space.
-The use of Chow-Liu trees was proposed by Buller and Kaufer [35] for estimating the entropy of multivariate data sources where the range of possible values is high.
-In our Python implementation, we use the pgmpy [46] library’s TreeSearch module.
-Practically, for each sensor subset, we:
+The tree captures dominant pairwise dependencie and significantly reduces computation time compared to naïve enumeration of the full joint measurement space.
+The use of Chow-Liu trees was proposed by Buller and Kaufer [37] for estimating the entropy of multivariate data sources where the range of possible values is high, as is the case with sensor outputs considered in this work.
+In our Python implementation, we use the pgmpy [53] library’s TreeSearch module to this end; practically, for each sensor subset, we:
   1. 1.
@@
   2. 2.
-    Build a Chow–Liu tree from the mutual information of each sensor pair, selecting edges to form a spanning tree.
+    Build a Chow–Liu tree from pairwise mutual information of sensors.
   3. 3.
-    Traverse the resulting tree to estimate max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$), and min-entropy ($H_{\infty}$) without enumerating the full exponential state space.
-Our framework evaluates the joint entropy over all sensor combinations.
+    Traverse the tree to estimate max ($H_{0}$), Shannon ($H_{1}$), collision ($H_{2}$), and min-entropy ($H_{\infty}$) without enumerating the full state space.
+Table 3: Average entropy results of Chow-Liu (CL) tree approximation against direct computation across all small sensor subsets ($n\in\{2,3\}$), giving the mean absolute error (MAE), averaged across all subsets, and the % mean relative error.
+$H_{0}$ $H_{1}$ $_{2} HH_{\infty}$
+Dataset $n$ Direct CL Direct CL Direct CL Direct CL MAE %
+UCI-HAR 2 17.60 16.92 12.08 12.58 10.85 10.33 6.01 6.20 0.473 4.1
+3 23.93 25.46 19.35 18.25 14.03 14.77 9.40 8.95 0.955 5.7
+Relay 2 15.36 15.54 9.72 10.12 8.27 7.95 4.64 4.88 0.285 3.0
+3 21.47 22.84 16.84 15.89 13.69 13.04 7.41 7.96 0.875 5.5
+PerilZIS 2 20.33 22.18 14.77 14.00 11.40 12.18 7.21 8.52 1.013 7.6
+3 30.82 32.55 21.28 20.08 16.66 17.65 11.46 12.06 1.258 6.2
+SHL 2 24.63 26.25 18.76 18.20 14.17 14.28 9.50 8.92 0.693 4.1
+3 36.71 39.05 28.44 26.83 22.46 21.14 12.44 13.25 1.488 4.9
+We validated the approach against a ground-truth derived by computing the joint entropy through exhaustive enumeration.
+As this direct method is only computationally feasible for a small number of modalities, we performed the comparison on all two- and three-sensor subsets from each dataset.
+The results of this validation are shown in Table 3.
+Across all datasets, the estimates from the Chow-Liu tree closely track direct ground-truth calculations in these low-dimensional cases.
+The relative error is consistently low (3–8%), supporting its use of Chow-Liu trees for higher-dimensional combinations where direct computation is infeasible.
+[^5] We note the caveat that efficiently estimating the entropy of high-dimensionality sources remains an ongoing research challenge [37].
+We must emphasise that our results are indicative estimations; quantifying errors beyond $n\geq 4$ is an open problem.
+Following this, we proceeded to estimating the entropy for all modalities in each dataset; our developed framework evaluates the joint entropy over all sensor combinations.
 The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module.
@@
 However, min-entropy ($H_{\infty}$) remains stubbornly low.
-For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ (Shannon) but saturates at only 21 bits of $H_{\infty}$.
+For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ but saturates at only 21 bits of $H_{\infty}$.
 Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all.
@@
 (d) SHL
-Figure 3: Entropy values versus the number of joint sensor modalities.
-[…]
+Figure 3: Average joint entropy vs. number of combined sensors modalities.
+We plot the number of combined modalities for each dataset in Figure 3, showing the average (mean) joint entropy for different entropy measures.
+The data illustrates that, while $H$ values are tightly coupled initially, they diverge significantly as the number of joint modalities increases.
+All datasets demonstrated a trend of worst-case $H_{\infty}$ being significantly lower than other entropy values, often barely increasing relative to best-case metrics.
+For SHL and PerilZIS, for example, $H_{\infty}$ increases by approximately only 1–2 bits per modality, with $H_{0}$$H_{2}$ increasing far more significantly ($\Delta H_{0}\approx$ 10 bits and $\Delta H_{1}\approx$ 5 bits for every added modality).
+In Tables 47, we give the differences of $H_{1}$ and $H_{\infty}$ for the top combinations as $H_{1}-H_{\infty}$, quantifying the sharp decreases between Shannon and min-entropy as the number of combined modalities increases.
+We also use # to denote the number of modalities used in each row.
 The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality.
-Combining signals increases the apparent capacity for unpredictability, but correlations between sensors means that the min-entropy from a large sensor ensemble is not be substantially higher than a reduced subset thereof.
-We note that standards implore the use of min-entropy as a safer, worst-case metric [26, 33].
-It measures how close the distribution is to collapsing around the single most probable outcome that an adversary will target first.
-To the best of our knowledge, the entropy collapse from highly correlated sensor modalities has not been elicited in existing work.
-We provide the full sensor combination results in our open-source repository.2
+Combining signals increases the apparent capacity for unpredictability, but redundancies between sensors means that the min-entropy from a large sensor ensemble is not substantially higher than a reduced subset thereof.
+This aligns with the emphasis of using min-entropy for worst-case guarantees in modern standards [25, 24].
+To our knowledge, the extent of entropy collapse under correlated mobile modalities has not been elicited at this scale.
+Full combination results are available in our repository (see Footnote 3).
 Table 4: Top 10 best-performing sensor combinations (UCI-HAR; in bits).
-Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
-All sensors 48.061 25.089 17.116 11.008
-(Acc.{x,y,z}, Gyro.{y,z}) 39.403 22.835 15.999 10.113
-(Acc.{x,y,z}, Gyro.{x,y}) 39.886 21.456 15.332 9.900
-(Acc.{x,y,z}, Gyro.{x,z}) 40.222 21.470 15.040 9.411
-(Acc.{y,z}, Gyro.{x,y,z}) 40.228 21.306 14.698 9.274
-(Acc.{x,z}, Gyro.{x,y,z}) 40.293 21.260 14.575 9.154
-(Acc.{x,y,z}, Gyro.y) 31.228 18.804 13.893 9.065
-(Acc.{x,y}, Gyro.{x,y,z}) 40.273 21.155 14.208 8.721
-(Acc.{x,y,z}, Gyro.z) 31.564 18.775 13.833 8.682
-(Acc.{y,z}, Gyro.{y,z}) 31.570 18.583 13.400 8.386
+Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
+All sensors 6 48.061 25.089 17.116 11.008 14.081
+(Acc.{x,y,z}, Gyro.{y,z}) 5 39.403 22.835 15.999 10.113 12.722
+(Acc.{x,y,z}, Gyro.{x,y}) 5 39.886 21.456 15.332 9.900 11.556
+(Acc.{x,y,z}, Gyro.{x,z}) 5 40.222 21.470 15.040 9.411 12.059
+(Acc.{y,z}, Gyro.{x,y,z}) 5 40.228 21.306 14.698 9.274 12.032
+(Acc.{x,z}, Gyro.{x,y,z}) 5 40.293 21.260 14.575 9.154 12.106
+(Acc.{x,y,z}, Gyro.y) 4 31.228 18.804 13.893 9.065 9.739
+(Acc.{x,y}, Gyro.{x,y,z}) 5 40.273 21.155 14.208 8.721 12.434
+(Acc.{x,y,z}, Gyro.z) 4 31.564 18.775 13.833 8.682 10.093
+(Acc.{y,z}, Gyro.{y,z}) 4 31.570 18.583 13.400 8.386 10.197
 Table 5: Top 10 best-performing sensor combinations (Relay; in bits).
-Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
-All sensors 45.794 25.036 15.725 8.092
-(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 44.795 24.963 15.334 7.859
-(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 38.385 21.590 14.783 7.702
-(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 37.026 21.001 14.553 7.702
-(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 36.092 20.794 14.450 7.673
-(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 37.385 21.517 14.428 7.469
-(Acc., Gyro., Light, Mag., Rot. Vec.) 36.026 20.924 14.270 7.465
-(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 35.092 20.720 14.147 7.439
-(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 41.094 22.794 14.437 7.338
-(Acc., Grav., Light, Mag., Rot. Vec.) 29.617 17.555 13.244 7.312
+Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
+All sensors 7 45.794 25.036 15.725 8.092 16.944
+(Acc., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 6 44.795 24.963 15.334 7.859 17.104
+(Acc., Grav., Light, Lin. Acc., Mag., Rot. Vec.) 6 38.385 21.590 14.783 7.702 13.888
+(Acc., Grav., Gyro., Light, Mag., Rot. Vec.) 6 37.026 21.001 14.553 7.702 13.299
+(Grav., Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 6 36.092 20.794 14.450 7.673 13.121
+(Acc., Light, Lin. Acc., Mag., Rot. Vec.) 5 37.385 21.517 14.428 7.469 14.048
+(Acc., Gyro., Light, Mag., Rot. Vec.) 5 36.026 20.924 14.270 7.465 13.459
+(Gyro., Light, Lin. Acc., Mag., Rot. Vec.) 5 35.092 20.720 14.147 7.439 13.281
+(Acc., Grav., Gyro., Light, Lin. Acc., Mag.) 6 41.094 22.794 14.437 7.338 15.456
+(Acc., Grav., Light, Mag., Rot. Vec.) 5 29.617 17.555 13.244 7.312 10.243
 Table 6: Top 10 best-performing sensor combinations (PerilZIS; in bits).
-Modalities $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
-All sensors 83.662 36.998 33.682 23.926
-(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 76.907 34.258 31.510 23.835
-(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.246 31.509 23.835
-(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.820 34.243 31.508 23.829
-(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 72.737 34.229 31.501 22.945
-(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 71.384 32.586 30.190 22.945
-(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 68.568 34.218 31.500 22.945
-(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 68.650 34.215 31.499 22.945
-(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 67.214 32.575 30.189 22.945
-(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 76.618 33.816 31.035 21.915
+Modalities # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
+All sensors 13 83.662 36.998 33.682 23.926 13.072
+(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 76.907 34.258 31.510 23.835 10.423
+(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.246 31.509 23.835 10.411
+(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.820 34.243 31.508 23.829 10.414
+(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.229 31.501 22.945 11.284
+(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 10 71.384 32.586 30.190 22.945 9.641
+(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.568 34.218 31.500 22.945 11.273
+(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.650 34.215 31.499 22.945 11.270
+(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 9 67.214 32.575 30.189 22.945 9.630
+(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 11 76.618 33.816 31.035 21.915 11.901
 Table 7: Top 10 best-performing sensor combinations (SHL; in bits).
-Modality $H_{0}$ $H_{1}$ $H_{2}$ $H_{\infty}$
-All sensors 158.601 82.301 39.320 21.289
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 148.995 78.624 39.276 21.289
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 148.759 78.477 39.272 21.289
-(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 148.587 78.380 39.249 21.289
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 148.952 78.135 39.235 21.289
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 139.153 74.800 39.175 21.289
-(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 138.981 74.703 39.128 21.289
-(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 138.745 74.556 39.117 21.289
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 139.346 74.458 39.100 21.289
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 139.109 74.311 39.087 21.289
+Modality # $H_{0}$ $H_{1}$ $_{2} HH_{\infty}$ $H_{1}-H_{\infty}$
+All sensors 22 158.601 82.301 39.320 21.289 61.012
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 148.995 78.624 39.276 21.289 57.335
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 148.759 78.477 39.272 21.289 57.188
+(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 21 148.587 78.380 39.249 21.289 57.091
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 20 148.952 78.135 39.235 21.289 56.846
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 19 139.153 74.800 39.175 21.289 53.511
+(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 138.981 74.703 39.128 21.289 53.414
+(Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 138.745 74.556 39.117 21.289 53.267
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 19 139.346 74.458 39.100 21.289 53.169
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 19 139.109 74.311 39.087 21.289 53.022
+4.3.4 Binning Sensitivity Analysis
+As we discussed in §4.1 and §4.3.1, bin width is a key consideration in estimating entropy accurately and efficiently.
+Too few bins can lead to oversmoothing, where distinct features of the data distribution are lost, artificially deflating the entropy estimate.
+Conversely, too many bins can lead to sparsity and overfitting, where the analysis models the specific sample’s noise rather than the source’s true underlying distribution, inflating entropy.
+We performed a sensitivity analysis on the best-performing sensor combination (“All sensors”) for each of our four datasets.
+We re-calculated the joint entropy while varying the number of bins for each modality, starting from 5 bins and increasing up to 2048.
+Beyond that, the computational cost of processing the large state space became prohibitive.
+The results of this analysis are presented in Figure 4.
+The graphs show a consistent trend across all datasets.
+For a low number of bins, all entropy measures are severely underestimated as expected.
+As the number of bins increases, the entropy rises sharply before tapering off and converging towards a stable plateau that approximates a logarithmic shape.
+(An exception is the min-entropy value for SHL, which converges after approximately 100 bins).
+Crucially, FD-selected bin counts fall within the stable region, supporting FD as a balanced choice between fidelity and tractability.
+(a) UCI-HAR
+(b) PerilZIS
+(c) Relay
+(d) SHL
+Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set. Freedman–Diaconis (FD) values are also shown.
 5 Evaluation
-This section critically evaluates our sensor entropy findings using existing literature and relevant standards.
-We then discuss potential mitigations such as randomness extractors, outline inherent limitations of our study and sensor-based approaches, and propose avenues for future enhancements.
+This section interprets our entropy findings for system design: we discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
 5.1 Discussion
-Prior studies have claimed or implied that mobile sensors are suitable data sources for security-critical applications.
+Prior studies have claimed that mobile sensors are suitable data sources for supporting security-critical applications.
 Much work has relied on model evaluation metrics as a proxy for evaluating such claims [7, 6, 21, 9], with a smaller subset using more established entropy metrics [36, 15].
-However, analyses have demonstrated that individual sensors confer very little entropy, especially under worst-case assumptions [17, 16, 15, 7].
-Our results, derived from an analysis of multiple datasets and modalities (see Tables 25), corroborate and challenge these claims.
+Separate entropy analyses of sensors have demonstrated that individual sensors actually confer very little entropy, particularly under worst-case assumptions [28, 27, 26, 17].
+Our results both challenge and extend these claims.
 Single sensors yield low entropy.
@@
 some single-sensor readings yield only 1–3 bits of min-entropy).
-This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not fully capture.
-Multiple sensors improve worst-case entropy, but not dramatically.
-While combining more modalities typically raises the upper bound (e.g.
-Hartley or Shannon entropy), our experiments reveal that min-entropy gains are far smaller than one might hope.
-In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution.
-Our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches.
-We estimate that even the most complicated multi-modal combination provide relatively little worst-case entropy.
-It is important to note that our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further.
-Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone source.
-They might be useful for tasks where strong assurances are not required, e.g.
-simple proximity checks; however, relying on sensor data as robust entropy sources is fraught with security risks.
-What can be done?
-One might hope that cryptographic extractors (e.g.
-Von Neumann extractors or more advanced schemes [47, 48, 49]) could make low-entropy sensors suitable for usage.
-Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of unpredictability beyond the source’s intrinsic min-entropy.
-If combined sensor data provides, say, 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more.
-This means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance).
-Extraction enhances the quality of the original source if it has enough min-entropy, but it cannot elevate a source that fundamentally lacks it.
-Another avenue is to introduce additional unpredictability through user interaction or deliberate environmental perturbation.
-For example, requiring the user to perform a random shaking gesture during a pairing protocol could inject extra entropy into the sensor readings.
-Prior studies have found that deliberate device movements can, indeed, increase the usable entropy of motion sensors, albeit only by a modest amount (on the order of a few bits) [15, 17].
-While this boost is non-negligible, it remains far below the dozens or hundreds of bits typically desired for security-critical applications.
-One possibility is to incorporate sensor-based randomness into an entropy pool that seeds a cryptographically secure pseudo-random number generator (CSPRNG).
-By doing so, a hybrid design can mitigate bias in the raw sensor data and produce outputs that pass statistical randomness tests (see the designs by Suciu et al.
-[27] and Wallace et al.
-[29]).
-However, this approach implicitly contradicts the rationale for many existing sensor-based applications, which leverage similarity or correlation in sensor readings.
-For example, in matching similar motion traces or environmental conditions of two neighbouring devices [42, 13, 2, 3, 4, 1].
-The desired property here is not pure unpredictability but rather shared information between the sensor signals collected by one or more devices.
-Using CSPRNG-derived signals would make it difficult to distinguish whether signals originated from co-located devices using similarity-based metrics.
-This exposes a fundamental security paradox in existing work.
-We have shown that sensors have insufficient entropy, even when using multiple ones simultaneously.
-Yet, methods to increase that entropy would compromise the usefulness of schemes that rely on the similarity of one or more sensor signals.
-We see no straightforward remedy for the entropy shortfall of mobile sensors.
-Extractors, user-assisted randomness, and entropy pools address diHoweverfferent problems: reducing bias, injecting small amounts of fresh entropy, or improving output distribution.
-Nevertheless, they do not compensate for fundamentally weak signals, which we have shown to be the case across a range of modalities.
-Sensor data and its derivatives may still be appropriate for low-security purposes.
-However, relying on it alone for security applications is highly inadvisable due to the serious risk of adversaries who exploit statistical biases to reduce the effective search space.
-5.2 Limitations
-Despite our detailed analysis, this work has some notable limitations.
-Firstly, while we analyse four large datasets, they do not fully capture contexts all possible contexts.
-Some work has suggested that performing dedicated movements (e.g.
-gestures) can increase the amount of usable entropy from motion sensors by 5–6 bits [17, 15].
+This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not capture at this scale.
+Combining multiple sensors offers modest gains to worst-case min-entropy.
+While using multimodal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect.
+In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes.
+Even the most complex multimodal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards.
+This also aligns with prior work showing that Shannon entropy can significantly overstate security [54, 55, 24, 25].
+Hence, while mobile sensors can augment other authentication or key-generation processes, they do not suffice by themselves.
+Using multiple sensors may still improve usability and model accuracy, e.g.
+for coarse co-location checks, but they should not be treated as secure, unpredictable sources.
+5.2 Translating Entropy Bounds Into Guessing Attacks
+We sketch how low-entropy sensor data translates into concrete attack cost.
+Let $Z$ be the discretised joint sensor vector used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy.
+Ordering guesses by decreasing probability, the success probability after $q$ guesses satisfies
+$\Pr[\text{success within }q]\;\leq\;q\cdot 2^{-H_{\infty}(Z)}.$ (10)
+The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ [55].
+Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching [7, 6, 9, 21, 35, 20, 19].
+Let us consider an abstract authentication function, $\mathrm{Auth}(z)\to\{0,1\}$, which accepts or rejects an input vector $z\in Z$.
+An adversary’s task becomes easier through a similarity-based scheme: they only need to generate a candidate $z^{\prime}$ that is ‘close enough’ to a high-probability vector $z$.
+This means the effective success probability for each guess is, in fact, higher than the bound in Eq.
+10 suggests.
+The best-performing combination in our study yielded $H_{\infty}\approx 24$ bits, corresponding to a search space of $2^{24}$ ($\approx$16.7 million) possibilities.
+An adversary may pre-compute the most likely vectors and submit them to $\mathrm{Auth}$, requiring only $2^{23}$ guesses on average.
+This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys [23].
+Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach [42].
+In the worst cases, such as the single-sensor readings where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring approximately $2^{7}=128$ guesses on average.
+While rate-limiting the authentication function can mitigate this risk, it does not solve the root problem.
+If each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$.
+Rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
+5.3 What can be done?
+Our results highlight a core tension in sensor-based security design.
+Proposals for pairing, co-location inference, or shared-secret agreement typically depend on the similarity of sensor signals across devices.
+Yet the very regularities that enable similarity seemingly stem from low-entropy structure.
+In effect, these systems trade cryptographic unpredictability for application-level utility.
+The open question is whether a scheme can ever achieve both meaningful entropy and the cross-device similarity it depends on.
+We close by offering several recommendations for researchers in this area.
+Evaluate the threat of adversarial samples
+A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors [49, 7, 9, 10, 21, 8, 5, 6].
+We urge the community to consider the resilience of proposals to adversarial samples, particularly when using ‘fuzzy’ matching methods such as supervised learning models, and to translate those estimates to concrete costs, such as expected guesses.
+Quantify effects of deliberate interaction and induced phenomena
+User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits [26, 28].
+For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors.
+A key gap is to quantify how much $\Delta H_{\infty}$ such interventions can actually deliver across modalities.
+Our expectation is that any gains remain modest, but this trade-off between added entropy and practical feasibility remains an under-explored area of research.
+Decouple sensor similarity from unpredictability
+Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs.
+The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see [38, 40]).
+Cryptographic extractors (e.g.
+Von Neumann and later constructions [56, 57, 58]) or decorrelation techniques (e.g.
+PCA) may be considered to convert low-quality sensor noise into higher-entropy sources.
+However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy.
+That is, if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more.
+Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost.
+Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
+5.4 Scope and Limitations
+While we analyse four large datasets, they do not fully capture contexts all possible contexts.
+Some work has suggested that performing dedicated movements, such as gestures, can increase the amount of usable entropy from motion sensors in the region of approximately 5–6 bits [28, 26].
 Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action.
-Secondly, our choice of Freedman-Diaconis binning and Chow-Liu trees is a pragmatic compromise.
-Smaller bin sizes tend to over-simplify the distribution and underestimate entropy, whereas larger bins can lead to computational blowup.
-As such, although our approach outperforms naïve joint entropy estimation, it is still an approximation.
-Thirdly, we largely focus on a data-centric view of entropy and do not examine other threats, such as sensor spoofing [3], hardware attacks [34] and cross-device correlation attacks (see [50] against wireless body area networks).
-It is possible that these threats may reduce the entropy of sensors even further.
+Conversely, we do not examine entropy-reducing threats, such as sensor spoofing [10], hardware attacks [44], or cross-device correlation attacks [45].
+It is concievable that these threats may reduce the entropy of sensors even further.
+[^6] By way of example, we could consider a voltage-based fault injection attack that causes nondeterministic errors in the output of individual sensors, affecting the entropy of their output values.
+We also estimate unconditional entropies such that all measurements are independent and identically distributed (i.i.d.).
+This is an optimistic assumption, as sensor data is inherently time-correlated; a sensor’s next reading is often highly dependent on its previous readings.
+For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g.
+min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see [59] for additional properties).
+Thus, the unconditional entropy values we calculate in this work act as an upper bound.
+We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial.
+Nevertheless, we posit that this would only serve to decrease entropy, strengthening our core conclusion that mobile sensors are insecure data sources for security applications.
 6 Conclusion
-This paper provides a comprehensive analysis of sensor-derived entropy across multiple datasets and modalities.
-Our results expose a tension between the perceived and actual strength of sensor data for security applications.
-Even in the best-performing sensor combinations, seemingly suitable results using one metric collapse to insecure levels when using standard worst-case metrics.
-Notably, modalities that yield ‘good’ max- or Shannon entropies, representing the best- and average-case unpredictability, have insecure worst-case min-entropies.
-Consequently, sensor modalities that may appear robust have biases that may enable adversaries to predict the most probable values with minimal effort.
-Our findings also challenge the prevailing orthodoxy that model evaluation metrics (e.g.
-accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals.
-The vulnerability of mobile sensors to biased distributions significantly undermine their effectiveness as reliable entropy sources.
-We also cast doubt on the use of sensors with respect to their non-stationary and lack of reproducibility.
-These issues collectively contradict the criteria articulated in frameworks such as NIST SP 800-90B, which emphasise noise-source stationarity and protection from external influence.
-The effectiveness of countermeasures remains an open research challenge.
-Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs in mobile device security.
+This paper presented the largest study to date on the entropy of mobile sensors across four public datasets (UCI-HAR, SHL, Relay, and PerilZIS) and 25 modalities.
+Our results expose a gap between the perceived and actual strength of sensor data for security: even when max- or Shannon entropies appear high, worst-case min-entropy often collapses to insecure levels.
+Modalities that appear strong under average-case metrics, which can mask inherent biases, are insecure under worst-case measures advocated in modern security standards.
+Consequently, this creates the opportunity for adversaries to guess the most probable values with minimal effort.
+Our findings also challenge the common practice that model evaluation metrics (e.g.
+accuracy or EER) suffice to demonstrate that sensor-based security applications are, indeed, secure.
+The vulnerability of mobile sensors to biased distributions and inter-sensor correlations significantly undermines their effectiveness as reliable sources of unpredictability.
+The effectiveness of countermeasures remains an open and pressing research challenge without compromising security or utility.
+Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, paving the way for safer and more robust designs.
 In future work, we consider that a dynamic analysis is important to assess the stationarity issues with sensor data, where entropy varies according to user behavior or environment.
-Moreover, a user study could yield empirical bounds on how finely humans can intentionally manipulate motion sensors, revealing more realistic limits to sensor-based randomness in real-world scenarios.
-For example, our binning decision was a statistical one, rather one that reflects human usage; it is possible that real-world usage may reduce the resolution of useful sensor data, thus reducing entropy.
+A user study could yield empirical bounds on how finely humans can intentionally manipulate motion sensors, revealing more realistic limits to sensor-based randomness in real-world scenarios.
+Our binning decision was a statistical one, rather than one that reflects human usage; it is possible that real-world usage may reduce the resolution of useful sensor data, reducing entropy even further.
 Overall, while sensor-based data can reach limited levels of entropy under favourable conditions, the road to making such sources systematically secure and robust is long.
-The key takeaway is that substantial work is needed before sensors can be considered appropriate for security-critical applications.
+The key takeaway is that substantial work is needed before mobile sensors can be considered appropriate for security-critical applications.
 Acknowledgments

### v9

@@
 Traditional key generation protocols [8, 18] have historically relied on shared motion patterns in order to establish secure communication between co-located devices.
-Meanwhile, other proposals have incorporated environmental phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
+Meanwhile, other proposals have incorporated environmental phenomena, such as magnetic field characteristics and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
 Despite these advances, a crucial assumption remains under-examined: that sensor data provides sufficient entropy to thwart adversarial inference.
-Many existing studies rely on heuristic assessments or model classsification metrics, e.g.
-[21, 7, 6, 9, 22, 19, 20], that do not address two basic questions: (1) how much entropy do sensors actually provide, and (2) to what extent do multimodal sensor combinations improve entropy?
-For the latter, multimodal sensing has been advocated to counter sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated.
-The security implications of low-entropy sensor data can be direct and severe.
+Many existing studies rely on heuristic assessments or model classification metrics, e.g.
+[21, 7, 6, 9, 22, 19, 20], which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy?
+For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated.
+The security implications of low-entropy sensor data can be both direct and severe.
 If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it.
 For such reasons, modern security standards provide important quantitative thresholds for what can be considered ‘secure’.
-For instance, NIST SP 800-57 [23] states that security levels below 80 bits are broken for long-term use, with 112 bits as the current minimum and 128 bits recommended for security keys.
-Standards that govern dedicated entropy sources are even stricter: NIST SP 800-90B [24] expects entropy sources used for cryptographic random number generators (RNGs) to approach full per-bit entropy, [^2] More specifically, $\approx$1 bit ($\epsilon\leq 2^{-32}$ per-bit) of min-entropy per output bit.
-while AIS 20/31 [25] requires an RNG’s internal state to possess over 240 and 250 bits of min-entropy and Shannon entropy respectively.
-As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls short of these standards.
+For instance, NIST SP 800-57 [23] states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys.
+Standards that govern random number generators (RNGs) are stricter still.
+NIST SP 800-90B [24] effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 [25] mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state.
+As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls well below these thresholds.
 In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—[26] (1 sensor), [27] (10), [17] (2), [28] (3)—across multiple public datasets.
@@
 Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets.
-Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically 40–75% below Shannon entropy for the same signals.
-In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be systematically enumerable.
+Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically $\approx$40–75% below Shannon entropy for the same signals.
+In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be enumerable by an adversary.
 Our contributions are as follows:
@@
 This raises fundamental concerns with respect to security applications that rely on signals from single and multiple mobile sensors.
-Consequently, we do not recommend relying wholly on mobile sensors as entropy sources for security-critical applications.
+Consequently, we do not recommend relying entirely on mobile sensors as entropy sources for security-critical applications.
 The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics.
@@
 To our knowledge, the extent of entropy collapse under correlated mobile modalities has not been elicited at this scale.
-Full combination results are available in our repository (see Footnote 3).
+Full combination results are available in our repository (see Footnote 2).
 Table 4: Top 10 best-performing sensor combinations (UCI-HAR; in bits).
@@
 All sensors 13 83.662 36.998 33.682 23.926 13.072
-(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 76.907 34.258 31.510 23.835 10.423
-(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.246 31.509 23.835 10.411
-(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.820 34.243 31.508 23.829 10.414
-(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 72.737 34.229 31.501 22.945 11.284
-(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 10 71.384 32.586 30.190 22.945 9.641
-(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.568 34.218 31.500 22.945 11.273
-(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 9 68.650 34.215 31.499 22.945 11.270
-(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 9 67.214 32.575 30.189 22.945 9.630
-(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 11 76.618 33.816 31.035 21.915 11.901
+(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 12 76.907 34.258 31.510 23.835 10.423
+(Acc.{y,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 72.737 34.246 31.509 23.835 10.411
+(Acc.{x,z}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 72.820 34.243 31.508 23.829 10.414
+(Acc.{x,y}, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 11 72.737 34.229 31.501 22.945 11.284
+(Acc.{x,y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 11 71.384 32.586 30.190 22.945 9.641
+(Acc.y, Light, Temp., Press., Mag.{x,y,z}, Gyro.{x,y,z}) 10 68.568 34.218 31.500 22.945 11.273
+(Acc.x, Light, Temp., Pres., Mag.{x,y,z}, Gyro.{x,y,z}) 10 68.650 34.215 31.499 22.945 11.270
+(Acc.{y,z}, Light, Temp., Mag.{x,y,z}, Gyro.{x,y,z}) 10 67.214 32.575 30.189 22.945 9.630
+(Acc.{x,y,z}, Light, Temp., Pres., Mag.{x,z}, Gyro.{x,y,z}, Hum.) 12 76.618 33.816 31.035 21.915 11.901
 Table 7: Top 10 best-performing sensor combinations (SHL; in bits).
@@
 All sensors 22 158.601 82.301 39.320 21.289 61.012
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 148.995 78.624 39.276 21.289 57.335
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 148.759 78.477 39.272 21.289 57.188
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 21 148.995 78.624 39.276 21.289 57.335
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 21 148.759 78.477 39.272 21.289 57.188
 (Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y,z}, Pres., Alt., Temp.) 21 148.587 78.380 39.249 21.289 57.091
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 20 148.952 78.135 39.235 21.289 56.846
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 19 139.153 74.800 39.175 21.289 53.511
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,y}, Pres., Alt., Temp.) 21 148.952 78.135 39.235 21.289 56.846
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{z}, Pres., Alt., Temp.) 20 139.153 74.800 39.175 21.289 53.511
 (Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y,z}, Pres., Alt., Temp.) 20 138.981 74.703 39.128 21.289 53.414
 (Acc.{x,y,z}, Gyro.{x,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x,z}, Pres., Alt., Temp.) 20 138.745 74.556 39.117 21.289 53.267
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 19 139.346 74.458 39.100 21.289 53.169
-(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 19 139.109 74.311 39.087 21.289 53.022
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{y}, Pres., Alt., Temp.) 20 139.346 74.458 39.100 21.289 53.169
+(Acc.{x,y,z}, Gyro.{x,y,z}, Mag.{x,y,z}, Ori.{w,x,y,z}, Grav.{x,y,z}, LinAcc.{x}, Pres., Alt., Temp.) 20 139.109 74.311 39.087 21.289 53.022
 4.3.4 Binning Sensitivity Analysis

### v10

@@
 Beyond activity recognition and context inference, e.g.
-[1, 2, 3, 4], a long line of work proposes using these signals for security-critical tasks.
-This includes cryptographic key generation, zero-interaction device pairing, and continuous authentication [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20].
-Traditional key generation protocols [8, 18] have historically relied on shared motion patterns in order to establish secure communication between co-located devices.
-Meanwhile, other proposals have incorporated environmental phenomena, such as magnetic field characteristics and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
-Despite these advances, a crucial assumption remains under-examined: that sensor data provides sufficient entropy to thwart adversarial inference.
-Many existing studies rely on heuristic assessments or model classification metrics, e.g.
-[21, 7, 6, 9, 22, 19, 20], which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy?
-For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated.
+gjoreski2018university , wang2019enabling , anguita2013public , yin2024systematic , a long line of work proposes using these signals for security-critical tasks, such as cryptographic key generation, zero-interaction device pairing, and continuous authentication gurulian2017effectiveness , gurulian2018good , mehrnezhad2015tap , mayrhofer2009shake , shrestha2014drone , shrestha2018sensor , shepherd2017applicability , shi2011senguard , miettinen2014conxsense , li2013unobservable , li2020t2pair , riva2012progressive , krhovjak2007sources , xu2021key , abuhamad2020sensor , mayrhofer2021adversary .
+Traditional key generation protocols mayrhofer2009shake , xu2021key have historically relied on shared motion patterns in order to establish secure communication between co-located devices.
+Meanwhile, other proposals have incorporated environmental phenomena, such as magnetic field characteristics and thermal fluctuations, to mitigate relay attacks shrestha2014drone , shrestha2018sensor , markantonakis2024using , gurulian2017effectiveness , shepherd2017applicability and to reduce the frequency of user authentication prompts krhovjak2007sources , riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable .
+Despite these advances, a crucial assumption remains under-examined: that sensor data is robust against adversaries that aim to forge or reconstruct such data through statistical inference.
+Many existing proposals rely on heuristic assessments or model classification metrics, e.g.
+markantonakis2024using , mehrnezhad2015tap , gurulian2018good , shrestha2014drone , truong2014comparing , abuhamad2020sensor , mayrhofer2021adversary , which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy?
+For the latter, multimodal sensing has been proposed as a means to counteract sensor-specific weaknesses shrestha2014drone , markantonakis2024using , truong2014comparing , mehrnezhad2015tap , but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated.
 The security implications of low-entropy sensor data can be both direct and severe.
 If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it.
-For such reasons, modern security standards provide important quantitative thresholds for what can be considered ‘secure’.
-For instance, NIST SP 800-57 [23] states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys.
+Modern standards provide important quantitative thresholds for what can be considered ‘secure’ for such reasons.
+For example, NIST SP 800-57 nistsp80057 states that security levels below 80 bits are considered insecure for long-term use, with 112 bits as the current minimum and 128 bits recommended for cryptographic keys.
 Standards that govern random number generators (RNGs) are stricter still.
-NIST SP 800-90B [24] effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 [25] mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state.
-As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls well below these thresholds.
-In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—[26] (1 sensor), [27] (10), [17] (2), [28] (3)—across multiple public datasets.
-Our results reveal structural limits.
+NIST SP 800-90B turan2018recommendation effectively requires $\approx$1 bit of min-entropy per output bit ($\epsilon\leq 2^{-32}$), while AIS 20/31 bsi2024ais31 mandates over 240 and 250 bits of min-entropy and Shannon entropy respectively for an RNG’s internal state.
+In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—voris2011accelerometers (1 sensor), hennebert2013entropy (10), krhovjak2007sources (2), lv2020analysis (3)—across multiple public datasets.
+Our results reveal widespread weaknesses, both corroborating and expanding on existing results.
 Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets.
-Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically $\approx$40–75% below Shannon entropy for the same signals.
+Combining multiple sensors confers only modest gains: we explore how inter-sensor correlations and statistical bias contribute to a divergence between average- and worst-case entropy, with min-entropy typically $\approx$40–75% below Shannon entropy for the same signals.
+Even when combining over 20 sensors simultaneously, the (joint) min-entropy remains below $\approx$22 bits, far below the thresholds considered secure by the aforementioned standards.
 In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be enumerable by an adversary.
@@
 2.1 Related Work
-Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context [29, 7, 11, 5, 21, 6, 9]; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret [8, 30, 18, 20]; and (3) continuous authentication, which aims to verify users passively during normal interaction [16, 12, 13, 14, 31, 32, 33, 34].
-Representative designs include tap-to-pair schemes using accelerometers [7], induced shared vibrations [6], and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues [9, 21].
-These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence [8, 30].
+Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context halevi2012secure , mehrnezhad2015tap , shepherd2017applicability , gurulian2017effectiveness , markantonakis2024using , gurulian2018good , shrestha2014drone ; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret mayrhofer2009shake , pan2018universense , xu2021key , mayrhofer2021adversary ; and (3) continuous authentication, which aims to verify users passively during normal interaction riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable , patel2016continuous , mekruksavanich2021deep , li2018sensor , micallef2015aren .
+Representative designs include tap-to-pair schemes using accelerometers mehrnezhad2015tap , induced shared vibrations gurulian2018good , and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues shrestha2014drone , markantonakis2024using .
+These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence mayrhofer2009shake , pan2018universense .
 Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based.
-Across the aforementioned application areas, effectiveness is usually inferred from discriminative performance.
+Across the above application areas, effectiveness is usually inferred from discriminative performance, i.e.
+the system’s performance in classifying illegitimate vs.
+legitimate samples.
 A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g.
 cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1.
-This template appears in numerous systems, including [7, 9, 21, 11, 16, 14, 13, 31, 32, 6].
+This template appears in numerous systems, including mehrnezhad2015tap , shrestha2014drone , markantonakis2024using , shepherd2017applicability , riva2012progressive , li2013unobservable , miettinen2014conxsense , patel2016continuous , mekruksavanich2021deep , gurulian2018good .
 For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al.
-[19], Alzubaidi and Kalita [35], and Mayrhofer and Sigg [20].
-While discriminative performance metrics are informative for usability and functionality, they do not quantify entropy of the underlying sources, which is the focus of our analysis.
+abuhamad2020sensor , Alzubaidi and Kalita alzubaidi2016authentication , and Mayrhofer and Sigg mayrhofer2021adversary .
+Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature.
+Specifically, T2Pair by Li et al.
+li2020t2pair reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al.
+wu2024t2pair reports 51–54 bits.
+We are aware of no work that applies min-entropy as a worst-case measure of unpredictability turan2018recommendation , bsi2024ais31 , buller2016estimating , alzubaidi2016authentication , xu2021key , mayrhofer2021adversary .
 Adjacent evidence also highlights practical limits under tight timing constraints.
@@
 In that setting, Markantonakis et al.
-[21] (building on [5, 11]) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues.
-Taken together, the literature demonstrates functional feasibility but leaves open the core question of how unpredictable the underlying measurements really are, especially under multimodal fusion.
-This motivates our entropy-centric evaluation in the remainder of the paper.
-Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature.
-Specifically, T2Pair by Li et al.
-[15] reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al.
-[36] reports 51–54 bits.
-Even in the best cases, these values are low relative to modern cryptographic standards; moreover, we are aware of no work using standard min-entropy as a worst-case measure of unpredictability [24, 25, 37] (see comparison tables of state-of-the-art schemes in [35, 18, 20]).
-We address this gap by analysing the entropy of a wide range of mobile sensors, both in isolation and in combination.
-Concurrent work has attempted to assess the entropy bestowed by mobile sensors as a separate line of inquiry.
+markantonakis2024using (building on gurulian2017effectiveness , shepherd2017applicability ) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues.
+This points to wider fragility issues with sensor-based security schemes in practice.
+Some studies have attempted to assess the entropy bestowed by mobile sensors on a single-sensor basis.
 Voris et al.
-[26] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone.
+voris2011accelerometers investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone.
 The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy.
-Intrinsic noise from the sensor’s circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation.
+Intrinsic noise from the sensor circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation.
 Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer.
 Lv et al.
-[28] analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone.
+lv2020analysis analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone.
 Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method.
 Krhovják et al.
-[17] examined image and audio data collected from mobile phone cameras and microphones respectively on a Nokia N73 and E-Ten X500 and M700 phones.
+krhovjak2007sources examined image and audio data collected from mobile phone cameras and microphones, respectively, on a Nokia N73 and E-Ten X500 and M700 phones.
 The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera).
 Hennebert et al.
-[27] analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
-A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g.
-accelerometer and vibration sensors, produced the highest entropy.
+hennebert2013entropy analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1.
+Min-entropy values of 0–7.85 are given, with motion sensors, e.g.
+accelerometer and vibration sensors, producing the highest entropy.
 Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices.
 Suciu et al.
-[38] proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests [39], but without precise entropy estimates.
+suciu2011unpredictable proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests rukhin2001nist , but without precise entropy estimates.
 Wallace et al.
-[40] explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices.
+wallace2016toward explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices.
 Unlike dedicated TRNGs (e.g.
-ring oscillators, Johnson–Nyquist thermal noise, quantum sources [41]), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources.
+ring oscillators, Johnson–Nyquist thermal noise, quantum sources hurley2020quantum ), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources.
 Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability.
-In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 [42, 43, 41, 24, 25].
+In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 mai2017guessability , uellenbeck2013quantifying , hurley2020quantum , turan2018recommendation , bsi2024ais31 .
 Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@
 Reference Sensorsa Metricsb Results (bits) Context / Use-Case
-Voris et al. [26] Acc. $H_{\infty}$ 3.1–11.4 True random number generation
-Hennebert et al. [27] 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
-Lv et al. [28] Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
-Krhovják et al. [17] Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
-Li et al. [15] Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
-Wu et al. [36] Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
-This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinatons, for security systems
+Voris et al. voris2011accelerometers Acc. $H_{\infty}$ 3.1–11.4 True random number generation
+Hennebert et al. hennebert2013entropy 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
+Lv et al. lv2020analysis Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
+Krhovják et al. krhovjak2007sources Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
+Li et al. li2020t2pair Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
+Wu et al. wu2024t2pair Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
+This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinations, for mobile device sensors
   • a
@@
 $H_{\alpha}(X)\,=\,\frac{1}{1-\alpha}\,\log\!\Biggl(\sum_{x\in\mathcal{X}}p(x)^{\alpha}\Biggr).$ (1)
-We draw attention to four special cases of $\alpha$ that are widely used in the literature.
+We use four cases of $\alpha$ that are widely used in the literature.
 First, the Hartley (Max) Entropy (Eq.
@@
 $H_{0}(X)\equiv\lim_{\alpha\to 0}H_{\alpha}(X)\,=\,\log\Bigl|\bigl\{\,x\in\mathcal{X}:p(x)>0\bigr\}\Bigr|.$ (2)
-Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ as $\alpha\to 1$.
+Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ at $\alpha\to 1$.
 $H_{1}(X)\equiv\lim_{\alpha\to 1}H_{\alpha}(X)\,=\,-\sum_{x\in\mathcal{X}}p(x)\,\log p(x).$ (3)
@@
 It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage.
-Min-entropy (Eq.
+To account for this, min-entropy (Eq.
 5) is used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes; it is the value in the limit $\alpha\to\infty$:
@@
 $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty.
-We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards [24, 25].
+We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards turan2018recommendation , bsi2024ais31 .
 We also rely on joint entropy for assessing the entropy of multiple random variables.
@@
 3.1 Threat Model and Assumptions
-We consider an adversary aiming to compromise sensor-based security schemes (e.g., key generation, proximity detection, and continuous authentication).
+We consider an adversary aiming to compromise sensor-based security schemes (e.g.
+key generation, proximity detection, and continuous authentication).
 The adversary may gather extensive statistical data about how smartphone sensors behave in everyday usage; for instance, from widely available open datasets.
@@
 As such, they may resort to exhaustive enumeration of the measurement space if the target source’s entropy is low enough.
-We exclude capabilities such as fault injection and other hardware attacks (see [44]) as well as side-channel attacks such as cross-device correlation attacks [45].
+We exclude capabilities such as fault injection and other hardware attacks (see shepherd2021physical ) as well as side-channel attacks such as cross-device correlation attacks dautov2019effects .
 While those could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard them as outside the scope of this work.
@@
 3.2 High-level Methodology
-In this work, we focus on estimating population-level (“global”) sensor distributions by pooling all users within a dataset.
+In this study, we focus on estimating population-level (“global”) sensor distributions by pooling all users within a dataset.
 This choice reflects typical deployments in which systems must serve previously unseen users (e.g.
@@
 We collect large-scale sensor readings from public datasets spanning diverse activities and device usage.
-Modalities include motion, environmental, and orientation sensors (dataset details appear at the end of this section).
+Modalities include motion, environmental, and orientation sensors (dataset details are given at the end of this section).
   2. 2.
@@
   5. 5.
-    Interpretation.
+    Evaluation.
 Finally, we interpret the resulting entropy measures, focusing on whether sensor outputs remain sufficiently unpredictable against an informed adversary.
 We compare single-sensor versus multi-sensor scenarios to verify if combining modalities truly alleviates biases or simply adds redundant data susceptible to predictability concerns.
-Throughout, we note the constraints of NIST SP 800-90B, SP 800-22 [24], and AIS 20/31 [25] when applied to multi-sensor streams [37, 28].
+Throughout, we note the constraints of NIST SP 800-90B, SP 800-22 turan2018recommendation , and AIS 20/31 bsi2024ais31 when applied to multi-sensor streams buller2016estimating , lv2020analysis .
 These frameworks target univariate sources rather than joint entropy of complex multivariate processes, which is the focus of our work.
 (For example, NIST SP 800-90B focusses on assessing entropy sources with reduced, i.e.
-8-bit, outputs [37, 24]).
+8-bit, outputs buller2016estimating , turan2018recommendation ).
 To scope suitable data, we searched IEEE DataPort, Google Scholar, Google Dataset Search, and GitHub.
-Several ostensibly “open” datasets were unavailable for download or had restrictive licensing [46, 47, 48].
+Several ostensibly “open” datasets were unavailable for download or had restrictive licensing Mahbub_Btas2016_UMDAA02 , stragapede2023behavepassdb , acien2021becaptcha .
 We selected four datasets offering diverse contexts, consistent sampling, and documented modalities:
   • 1.
-    UCI-HAR [3]: A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities.
+    UCI-HAR anguita2013public : A widely referenced dataset for human activity recognition, comprising smartphone sensor recordings from multiple subjects performing daily activities.
 Data includes triaxial accelerometer and gyroscope signals.
   • 2.
-    University of Sussex–Huawei Locomotion (SHL) [1, 2]: Sampled at 100 Hz from an Huawei Mate 9 smartphone.
-TWe use the publicly available SHL Preview dataset (three recording days per user; 59 hours total) and, for scope alignment with prior work, the handheld-phone subset.
+    University of Sussex–Huawei Locomotion (SHL) gjoreski2018university , wang2019enabling : Sampled at 100 Hz from an Huawei Mate 9 smartphone.
+We use the publicly available SHL Preview dataset containing three recording days per user (59 hours total).
+For scope alignment with prior work, the handheld-phone subset is used.
   • 3.
-    Relay [5]: Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g.
-cafés), including accelerometer, gyroscope, and environmental readings in emulated payment scenarios.
+    Relay gurulian2017effectiveness : Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g.
+cafés), including accelerometer, gyroscope, and environmental readings taken from emulated payment scenarios.
   • 4.
-    PerilZIS [49]: Collected at 10 Hz from a TI SensorTag, Samsung Galaxy S6, and Samsung Galaxy Gear, spanning multiple zero-interaction security use cases in an office environment.
+    PerilZIS fomichev2019perils : Measurements collected at 10 Hz from a TI SensorTag, Samsung Galaxy S6, and Samsung Galaxy Gear, spanning multiple zero-interaction security use cases in an office environment.
 These datasets provide varied sensor types, user activities, and sampling rates, enabling us to observe how biases and inter-sensor correlations manifest across scenarios.
@@
 A crucial, yet underexplored, issue in prior work (e.g.
-[26, 28, 27, 42]) is how to appropriately estimate entropy from quantised sensor outputs.
-Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution.
-Yet, a sensor’s advertised resolution—for instance, 12 bits in the widely used Bosch BMA400 mobile accelerometer [50]—does not imply uniform histograms across its range of values.
+voris2011accelerometers , lv2020analysis , hennebert2013entropy , mai2017guessability ) is how to appropriately estimate entropy from quantised sensor outputs.
+Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with finite resolutions.
+Notably, a sensor’s advertised resolution—for instance, 12 bits in the widely used Bosch BMA400 mobile accelerometer bosch_bma400 —does not imply uniform histograms across its range of values.
 Everyday usage produces highly non-uniform biases and clustering, resulting in some measurements occurring more frequently than others.
@@
 triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al.
-[26].
+voris2011accelerometers .
 This approach allows for a fine-grained analysis of directional biases, as movement along one axis may be more constrained or predictable than others in typical usage scenarios.
@@
 This gap reflects distributions where a few outcomes dominate, driving worst-case unpredictability down even if the average-case picture is more favorable.
-Overall, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own.
+In general, the results confirm that data from individual sensors do not provide sufficient min-entropy for robust security on their own.
 In the next section, we examine whether combining multiple modalities can meaningfully increase this worst-case unpredictability or whether correlated biases persist across different sensor streams.
 4.3 Multi-modal Sensors
-Several sensor-based security proposals [7, 22, 21, 9, 10] assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than one.
-Let us now explore that claim.
-Initially, a naïve approach might add Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
+Several sensor-based security proposals mehrnezhad2015tap , truong2014comparing , markantonakis2024using , shrestha2014drone , shrestha2018sensor assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than a single one.
+An initial, naïve approach may attempt to model this by adding the Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
 However, this requires that $X_{i}$ are statistically independent.
-In reality, mobile sensors often exhibit strong dependencies; for instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices [51].
+In reality, mobile sensors often exhibit strong dependencies; for instance, the rotation vector, gravity, and linear acceleration sensors are frequently derived in software from the accelerometer and gyroscope on consumer devices android_motion_sensors .
 As a result, these modalities cannot be treated as independent variables.
@@
 Figure 2 provides a quantitative view of inter-sensor redundancy through Pearson correlation matrices for each dataset.
-Strong correlations, with $|\rho|\approx 1$, indicate high linear dependence.
-Such redundancy is a major contributor to fragility in pairwise combinations.
-Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and provides diminishing returns on min-entropy.
-In other words, while individual modalities may appear moderately unpredictable, the correlations shown in Figure 2 can sharply constrain the overall joint entropy compared to what would be expected if they were independent variables.
+Strong correlations, with $|\rho|\approx 1$, indicate high linear dependence; such redundancy is a major contributor to fragility in the context of pairwise combinations.
+Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and can provide diminishing returns.
+Consequently, while modalities may appear as offering some level of unpredictability individually, the correlations shown in Figure 2 can sharply constrain the overall joint entropy when compared to what would be expected if they were treated as independent variables.
 4.3.1 Complexity Challenges
@@
 Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset.
-Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates, as we discussed §4.1.
+Reducing bin sizes can help, but this risks oversimplifying the distribution and artificially deflating entropy estimates (discussed in §4.1).
 In our tests, we found that reducing bin counts reduced single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 2.
@@
 4.3.2 Chow-Liu Approximation
-To handle these scaling issues, we adopt Chow–Liu trees [52] for approximating a high-dimensional joint distribution by a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
+To handle these scaling issues, we adopt Chow–Liu trees chow1968approximating for approximating a high-dimensional joint distribution using a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
 Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies.
@@
 Chow–Liu trees are acyclic, singly connected structures: each node has at most one parent where one can traverse the tree to accumulate probabilities between pairwise dependencies.
-The tree captures dominant pairwise dependencie and significantly reduces computation time compared to naïve enumeration of the full joint measurement space.
-The use of Chow-Liu trees was proposed by Buller and Kaufer [37] for estimating the entropy of multivariate data sources where the range of possible values is high, as is the case with sensor outputs considered in this work.
-In our Python implementation, we use the pgmpy [53] library’s TreeSearch module to this end; practically, for each sensor subset, we:
+The tree captures dominant pairwise dependencies and significantly reduces computation time compared to naïve enumeration of the full joint measurement space.
+The use of Chow-Liu trees was proposed by Buller and Kaufer buller2016estimating for estimating the entropy of multivariate data sources where the range of possible values is high, as is the case with sensor outputs considered in this work.
+In our Python implementation, we use the pgmpy ankan2024pgmpy library’s TreeSearch module to this end; practically, for each sensor subset, we:
   1. 1.
@@
 The results of this validation are shown in Table 3.
-Across all datasets, the estimates from the Chow-Liu tree closely track direct ground-truth calculations in these low-dimensional cases.
-The relative error is consistently low (3–8%), supporting its use of Chow-Liu trees for higher-dimensional combinations where direct computation is infeasible.
-[^4] We note the caveat that efficiently estimating the entropy of high-dimensionality sources remains an ongoing research challenge [37].
+Across all datasets, the estimates from Chow-Liu trees closely track direct ground-truth calculations in these low-dimensional cases.
+The relative error is consistently low (3–8%), supporting its use within higher-dimensional combinations where direct computation is infeasible.
+[^4] We note the caveat that efficiently estimating the entropy of high-dimensionality sources remains an ongoing research challenge buller2016estimating .
 We must emphasise that our results are indicative estimations; quantifying errors beyond $n\geq 4$ is an open problem.
-Following this, we proceeded to estimating the entropy for all modalities in each dataset; our developed framework evaluates the joint entropy over all sensor combinations.
-The powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module.
-Processing all datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) and 32 GB RAM on Ubuntu 24.04.
+Following this, we proceeded to estimate the entropy for all modalities in each dataset.
+Our developed framework evaluates the joint entropy over all sensor combinations; the powerset of the sensor set is generated and processed in parallel using Python’s multiprocessing module.
+Processing all datasets took approximately 22 hours on our workstation with an Intel i7-6700K (8M cache, 4.20 GHz) CPU and 32 GB RAM on Ubuntu 24.04.
 4.3.3 Results
@@
 For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ but saturates at only 21 bits of $H_{\infty}$.
-Interestingly, we find that omitting certain correlated sensors sometimes does not reduce min-entropy at all.
+Interestingly, we find that omitting certain correlated sensors sometimes barely reduces min-entropy at all.
 For the Relay dataset in Table 5, the combination (Acc., Gyro., Light, Lin.
 Acc., Mag., Rot.
-Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the full set’s $8.092$ bits.
+Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the $8.092$ bits produced by the full of set of sensors.
 Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$.
@@
 The data illustrates that, while $H$ values are tightly coupled initially, they diverge significantly as the number of joint modalities increases.
-All datasets demonstrated a trend of worst-case $H_{\infty}$ being significantly lower than other entropy values, often barely increasing relative to best-case metrics.
+All datasets demonstrated a trend of worst-case $H_{\infty}$ remaining significantly lower and growing only marginally with respect to other entropy values.
 For SHL and PerilZIS, for example, $H_{\infty}$ increases by approximately only 1–2 bits per modality, with $H_{0}$$H_{2}$ increasing far more significantly ($\Delta H_{0}\approx$ 10 bits and $\Delta H_{1}\approx$ 5 bits for every added modality).
@@
 We also use # to denote the number of modalities used in each row.
-The results suggest that many sensors contribute redundant information when combined, exposing a fundamental limitation of multi-modal data in reality.
-Combining signals increases the apparent capacity for unpredictability, but redundancies between sensors means that the min-entropy from a large sensor ensemble is not substantially higher than a reduced subset thereof.
-This aligns with the emphasis of using min-entropy for worst-case guarantees in modern standards [25, 24].
+These results confirm that many sensors contribute little beyond redundancy when combined, revealing a fundamental limitation of multi-modal data in practice.
+While joint signals raise the apparent entropy, correlations between modalities prevent a corresponding rise in min-entropy; large ensembles offer no substantial advantage over smaller subsets of sensors.
+This underscores why min-entropy is the relevant metric for worst-case guarantees in modern standards bsi2024ais31 , turan2018recommendation .
 To our knowledge, the extent of entropy collapse under correlated mobile modalities has not been elicited at this scale.
@@
 As we discussed in §4.1 and §4.3.1, bin width is a key consideration in estimating entropy accurately and efficiently.
-Too few bins can lead to oversmoothing, where distinct features of the data distribution are lost, artificially deflating the entropy estimate.
-Conversely, too many bins can lead to sparsity and overfitting, where the analysis models the specific sample’s noise rather than the source’s true underlying distribution, inflating entropy.
+Too few bins cause oversmoothing, erasing distributional features and deflating entropy.
+At the same time, too many bins induces sparsity and overfitting, inflating entropy by modelling noise rather than the source.
 We performed a sensitivity analysis on the best-performing sensor combination (“All sensors”) for each of our four datasets.
@@
 5.1 Discussion
-Prior studies have claimed that mobile sensors are suitable data sources for supporting security-critical applications.
-Much work has relied on model evaluation metrics as a proxy for evaluating such claims [7, 6, 21, 9], with a smaller subset using more established entropy metrics [36, 15].
-Separate entropy analyses of sensors have demonstrated that individual sensors actually confer very little entropy, particularly under worst-case assumptions [28, 27, 26, 17].
+Prior studies have directly claimed or implicitly assumed that mobile sensors are suitable data sources for security-critical applications.
+To this end, work has relied on model evaluation metrics as a proxy for evaluating such claims mehrnezhad2015tap , gurulian2018good , markantonakis2024using , shrestha2014drone , with a smaller subset using more established entropy metrics wu2024t2pair , li2020t2pair .
+Separate entropy analyses of sensors have demonstrated that individual sensors actually confer very little entropy, particularly under worst-case assumptions lv2020analysis , hennebert2013entropy , voris2011accelerometers , krhovjak2007sources .
 Our results both challenge and extend these claims.
 Single sensors yield low entropy.
-Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g., some single-sensor readings yield only 1–3 bits).
+Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g.
+some single-sensor readings yield only 1–3 bits).
 Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g.
@@
 This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not capture at this scale.
-Combining multiple sensors offers modest gains to worst-case min-entropy.
-While using multimodal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect.
+Combining multiple sensors offers modest gains with respect to worst-case min-entropy.
+While using multi-modal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect.
 In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes.
-Even the most complex multimodal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards.
-This also aligns with prior work showing that Shannon entropy can significantly overstate security [54, 55, 24, 25].
+Even the most complex multi-modal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards.
+This also aligns with prior work showing that Shannon entropy can significantly overstate security arikan2002inequality , cachin1997entropy , turan2018recommendation , bsi2024ais31 .
 Hence, while mobile sensors can augment other authentication or key-generation processes, they do not suffice by themselves.
@@
 5.2 Translating Entropy Bounds Into Guessing Attacks
-We sketch how low-entropy sensor data translates into concrete attack cost.
+We now sketch how low-entropy sensor data translates into concrete attack cost.
 Let $Z$ be the discretised joint sensor vector used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy.
@@
 $\Pr[\text{success within }q]\;\leq\;q\cdot 2^{-H_{\infty}(Z)}.$ (10)
-The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ [55].
-Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching [7, 6, 9, 21, 35, 20, 19].
+The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ cachin1997entropy .
+Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching mehrnezhad2015tap , gurulian2018good , shrestha2014drone , markantonakis2024using , alzubaidi2016authentication , mayrhofer2021adversary , abuhamad2020sensor .
 Let us consider an abstract authentication function, $\mathrm{Auth}(z)\to\{0,1\}$, which accepts or rejects an input vector $z\in Z$.
@@
 An adversary may pre-compute the most likely vectors and submit them to $\mathrm{Auth}$, requiring only $2^{23}$ guesses on average.
-This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys [23].
-Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach [42].
+This is well within the capabilities of a moderately resourced adversary—taking seconds to milliseconds on a modern PC to enumerate—and is orders of magnitude weaker than the $2^{128}$ expected of standard cryptographic keys nistsp80057 .
+Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach mai2017guessability .
 In the worst cases, such as the single-sensor readings where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring approximately $2^{7}=128$ guesses on average.
-While rate-limiting the authentication function can mitigate this risk, it does not solve the root problem.
+A main defence against this is to employ rate-limit the number of attempts, although it does not solve the root problem.
 If each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$.
-Rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
+That is, rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
 5.3 What can be done?
@@
 Evaluate the threat of adversarial samples
-A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors [49, 7, 9, 10, 21, 8, 5, 6].
+A persistent feature in the literature is the lack of rigorous analysis of min-entropy (and induced guesswork) under realistic acceptance regions and attacker priors fomichev2019perils , mehrnezhad2015tap , shrestha2014drone , shrestha2018sensor , markantonakis2024using , mayrhofer2009shake , gurulian2017effectiveness , gurulian2018good .
 We urge the community to consider the resilience of proposals to adversarial samples, particularly when using ‘fuzzy’ matching methods such as supervised learning models, and to translate those estimates to concrete costs, such as expected guesses.
 Quantify effects of deliberate interaction and induced phenomena
-User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits [26, 28].
+User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits voris2011accelerometers , lv2020analysis .
 For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors.
@@
 Where applications need similarity evidence, such as for device authentication, one should treat that evidence as weak, predictable inputs.
-The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see [38, 40]).
+The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see suciu2011unpredictable , wallace2016toward ).
 Cryptographic extractors (e.g.
-Von Neumann and later constructions [56, 57, 58]) or decorrelation techniques (e.g.
+Von Neumann and later constructions von195113 , trevisan2001extractors , barak2003true ) or decorrelation techniques (e.g.
 PCA) may be considered to convert low-quality sensor noise into higher-entropy sources.
-However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy.
-That is, if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more.
+However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy; if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more.
 Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost.
@@
 While we analyse four large datasets, they do not fully capture contexts all possible contexts.
-Some work has suggested that performing dedicated movements, such as gestures, can increase the amount of usable entropy from motion sensors in the region of approximately 5–6 bits [28, 26].
+Some work has suggested that performing dedicated movements, such as gestures, can increase the amount of usable entropy from motion sensors in the region of approximately 5–6 bits lv2020analysis , voris2011accelerometers .
 Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action.
-Conversely, we do not examine entropy-reducing threats, such as sensor spoofing [10], hardware attacks [44], or cross-device correlation attacks [45].
+Conversely, we do not examine entropy-reducing threats, such as sensor spoofing shrestha2018sensor , hardware attacks shepherd2021physical , or cross-device correlation attacks dautov2019effects .
 It is concievable that these threats may reduce the entropy of sensors even further.
@@
 For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g.
-min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see [59] for additional properties).
+min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see iwamoto2013information for additional properties).
 Thus, the unconditional entropy values we calculate in this work act as an upper bound.
@@
 Our findings also challenge the common practice that model evaluation metrics (e.g.
-accuracy or EER) suffice to demonstrate that sensor-based security applications are, indeed, secure.
+accuracy or EER) suffice to demonstrate that sensor-based security applications are, indeed, secure from a data-centric perspective.
 The vulnerability of mobile sensors to biased distributions and inter-sensor correlations significantly undermines their effectiveness as reliable sources of unpredictability.

### v12

@@
 Using direct computation and estimation, we report entropy values—max, Shannon, collision, and min-entropy—for an exhaustive range of sensor combinations.
-We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, such as for authentication and key generation systems, even when many sensors are combined.
+We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, even when many sensors are combined.
 In particular, we observe an alarming divergence between average-case Shannon entropy and worst-case min-entropy.
-Single-sensor mean min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher.
-We further show that correlations between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy.
-This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits when combining 20+ sensor modalities.
+Single-sensor min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher.
+We also show that redundancies between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy.
+Indeed, min-entropy plateaus between 8.1–23.9 bits when combining up to 22 modalities, while Shannon entropy can exceed 80 bits.
+Adding sensors typically increases Shannon entropy but moves min-entropy by only $\approx$1–2 bits per added modality, evidencing entropy collapse under redundancy.
 Our results reveal that adversaries may feasibly predict sensor signals through an exhaustive exploration of the measurement space.
@@
 Meanwhile, other proposals have incorporated environmental phenomena, such as magnetic field characteristics and thermal fluctuations, to mitigate relay attacks shrestha2014drone , shrestha2018sensor , markantonakis2024using , gurulian2017effectiveness , shepherd2017applicability and to reduce the frequency of user authentication prompts krhovjak2007sources , riva2012progressive , shi2011senguard , miettinen2014conxsense , li2013unobservable .
-Despite these advances, a crucial assumption remains under-examined: that sensor data is robust against adversaries that aim to forge or reconstruct such data through statistical inference.
-Many existing proposals rely on heuristic assessments or model classification metrics, e.g.
+Despite these advances, a crucial assumption remains under-examined: that sensor data is robust against adversaries aiming to forge or reconstruct such data through statistical inference.
+Existing proposals rely heavily on heuristic assessments or model classification metrics, e.g.
 markantonakis2024using , mehrnezhad2015tap , gurulian2018good , shrestha2014drone , truong2014comparing , abuhamad2020sensor , mayrhofer2021adversary , which do not address two basic questions: (1) how much entropy do sensors actually provide and (2) to what extent do multimodal sensor combinations improve entropy?
@@
 The security implications of low-entropy sensor data can be both direct and severe.
-If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it.
+If such data is used as authentication evidence, or to derive secret keys, an adversary may attempt to enumerate the entire space of possible sensor measurement values in order to recover authentication or key material.
 Modern standards provide important quantitative thresholds for what can be considered ‘secure’ for such reasons.
@@
 Representative designs include tap-to-pair schemes using accelerometers mehrnezhad2015tap , induced shared vibrations gurulian2018good , and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues shrestha2014drone , markantonakis2024using .
-These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence mayrhofer2009shake , pan2018universense .
-Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based.
+Proposals commonly combine sensor modalities to gather additional evidence of co-presence; for example, using accelerometer and gyroscope measurements to capture shared device motion in addition to environmental readings mayrhofer2009shake , pan2018universense , alzubaidi2016authentication , xu2021key .
+A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g.
+cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1.
+This template appears in numerous systems, including mehrnezhad2015tap , shrestha2014drone , markantonakis2024using , shepherd2017applicability , riva2012progressive , li2013unobservable , miettinen2014conxsense , patel2016continuous , mekruksavanich2021deep , gurulian2018good .
 Across the above application areas, effectiveness is usually inferred from discriminative performance, i.e.
@@
 legitimate samples.
-A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g.
-cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1.
-This template appears in numerous systems, including mehrnezhad2015tap , shrestha2014drone , markantonakis2024using , shepherd2017applicability , riva2012progressive , li2013unobservable , miettinen2014conxsense , patel2016continuous , mekruksavanich2021deep , gurulian2018good .
-For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al.
-abuhamad2020sensor , Alzubaidi and Kalita alzubaidi2016authentication , and Mayrhofer and Sigg mayrhofer2021adversary .
+(For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al.
+abuhamad2020sensor , Alzubaidi and Kalita alzubaidi2016authentication , and Mayrhofer and Sigg mayrhofer2021adversary ).
+Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based.
 Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature.
@@
 This points to wider fragility issues with sensor-based security schemes in practice.
-Some studies have attempted to assess the entropy bestowed by mobile sensors on a single-sensor basis.
+Some studies have attempted to measure the entropy of single mobile sensors.
 Voris et al.
-voris2011accelerometers investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone.
+voris2011accelerometers evaluated accelerometers as true RNGs (TRNGs) on a WISP RFID tag and Nokia N97 phone.
 The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy.
@@
 accelerometer and vibration sensors, producing the highest entropy.
-Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices.
+Sensors have also been suggested as entropy sources within low-cost RNG designs for mobile devices.
 Suciu et al.
@@
 In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 mai2017guessability , uellenbeck2013quantifying , hurley2020quantum , turan2018recommendation , bsi2024ais31 .
+While several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts.
+Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges.
 Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@
 Reference Sensorsa Metricsb Results (bits) Context / Use-Case
-Voris et al. voris2011accelerometers Acc. $H_{\infty}$ 3.1–11.4 True random number generation
+Voris et al. voris2011accelerometers Acc $H_{\infty}$ 3.1–11.4 True random number generation
 Hennebert et al. hennebert2013entropy 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
-Lv et al. lv2020analysis Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
-Krhovják et al. krhovjak2007sources Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
-Li et al. li2020t2pair Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
-Wu et al. wu2024t2pair Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
+Lv et al. lv2020analysis Acc, Gyro, Mag $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
+Krhovják et al. krhovjak2007sources Cam, Mic $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
+Li et al. li2020t2pair Mic, Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
+Wu et al. wu2024t2pair Mic, Light $H_{1}$ 51–54 Zero-interaction device pairing
 This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinations, for mobile device sensors
   • a
-    Acc.: Accelerometer; Gyr.: Gyroscope; Mag.: Magnetometer; Cam.: Camera; Mic.: Microphone.
+    Acc: Accelerometer; Gyro: Gyroscope; Mag: Magnetometer; Cam: Camera; Mic: Microphone.
   • b
     $H_{0}$: Max; $H_{1}$: Shannon; $H_{2}$: Collision; $H_{\infty}$: Min-entropy.
-In general, while several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts.
-Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges.
 2.2 Definitions
@@
 We collect large-scale sensor readings from public datasets spanning diverse activities and device usage.
-Modalities include motion, environmental, and orientation sensors (dataset details are given at the end of this section).
+Modalities include motion, environmental, and orientation sensors (details are given at the end of this section).
   2. 2.
@@
   • 3.
-    Relay gurulian2017effectiveness : Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g.
+    Relay markantonakis2024using : Approximately 1,500 NFC-based contactless transactions recorded at 100 Hz across several physical locations (e.g.
 cafés), including accelerometer, gyroscope, and environmental readings taken from emulated payment scenarios.
@@
 These datasets provide varied sensor types, user activities, and sampling rates, enabling us to observe how biases and inter-sensor correlations manifest across scenarios.
-Next, we describe preprocessing and aggregation to form the global distributions used in our entropy analyses.
+We summarise the datasets in Table 2.
+Table 2: Summary of datasets.
+Dataset Devices Sensorsa Freq. Contextb
+UCI-HAR anguita2013public Samsung Galaxy S2 Acc, Gyro 50Hz ADL; 30 users
+SHL gjoreski2018university , wang2019enabling Huawei Mate 9 Acc, Gyro, Mag, Grav, LinAcc, RotVec, Light, Temp, Pres, Alt 100Hz HAR; transport; 3 users; 59 hrs
+Relay markantonakis2024using Google Nexus 5, 9; Samsung Galaxy S4, SGS5 Acc, Gyro, Mag, RotVec, Grav, LinAcc, Light 100Hz NFC transactions; 252 users
+PerilZIS fomichev2019perils TI SensorTag; Samsung Galaxy S6, Gear Acc, Gyro, Mag, Light, Temp, Pres 10Hz ZIS; 4 users; 8 devices; 4250 hrs
+  • a
+    Acc: Accelerometer; Gyro: Gyroscope; Mag: Magnetometer; Grav: Gravity; LinAcc: Linear acceleration; RotVec: Rotation vector; Temp: Temperature; Pres: Pressure; Alt: Altitude.
+  • b
+    ADL: Activities of daily living; HAR: Human activity recognition; NFC: Near-field communication; ZIS: Zero-interaction security.
 4 Entropy Analysis
@@
 4.3 Multi-modal Sensors
-Several sensor-based security proposals mehrnezhad2015tap , truong2014comparing , markantonakis2024using , shrestha2014drone , shrestha2018sensor assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than a single one.
+Several sensor-based proposals mehrnezhad2015tap , truong2014comparing , markantonakis2024using , shrestha2014drone , shrestha2018sensor assert that combining multiple sensor modalities bolsters security, on the intuition that an adversary must accurately predict several data streams rather than a single one.
 An initial, naïve approach may attempt to model this by adding the Shannon entropies from individual sensors, benefitting from the relation $H(X_{1},\ldots,X_{n})=\sum_{i=1}^{n}H(X_{i})$.
@@
 Employing correlated sensors adds little new unpredictability, which invalidates the simplistic additive model of entropy and can provide diminishing returns.
-Consequently, while modalities may appear as offering some level of unpredictability individually, the correlations shown in Figure 2 can sharply constrain the overall joint entropy when compared to what would be expected if they were treated as independent variables.
+Figure A.1 (Appendix A) presents mutual information heatmaps.
+These results reveal dependencies between sensor pairs, confirming that redundancy extends beyond what Pearson correlation alone suggests.
+Consequently, while modalities may appear as offering some level of unpredictability individually, the results suggest that overall joint entropy can be sharply constrained compared to what would be expected if they were treated as independent variables.
 4.3.1 Complexity Challenges
@@
 4.3.2 Chow-Liu Approximation
-To handle these scaling issues, we adopt Chow–Liu trees chow1968approximating for approximating a high-dimensional joint distribution using a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
+To handle these scaling issues, we adopt Chow–Liu trees chow1968approximating to approximate a high-dimensional joint distribution using a maximum-weight spanning tree, $\pi$, over the different sensor modalities.
 Each edge is weighted by the mutual information of the connected variables, ensuring the tree structure captures the dominant pairwise dependencies.
 This approach minimises the Kullback–Leibler divergence (Def.
-2.2) between the true multivariate distribution and the resulting tree-based approximation as follows:
+2.2) between the true multivariate distribution and the tree-based approximation as follows:
 $p_{\pi}(x_{1},\dots,x_{n})=p(x_{r})\prod_{i\neq r}p(x_{i}\;|\;x_{\pi(x)})$ (9)
@@
 Tables 58 report the top 10 performing multi-sensor combinations ranked by min-entropy for each dataset.
-As expected, combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy.
-However, min-entropy ($H_{\infty}$) remains stubbornly low.
-For instance, the complete set of sensors in SHL surpasses 80 bits of $H_{1}$ but saturates at only 21 bits of $H_{\infty}$.
+Combining all sensors yields the highest $H_{0}$ (max-entropy) and often increases Shannon and collision entropy.
+However, min-entropy ($H_{\infty}$) remains stubbornly low; the complete set of sensors in SHL, for instance, surpasses 80 bits of $H_{1}$, but saturates at only 21 bits of $H_{\infty}$.
 Interestingly, we find that omitting certain correlated sensors sometimes barely reduces min-entropy at all.
@@
 Acc., Mag., Rot.
-Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the $8.092$ bits produced by the full of set of sensors.
+Vec.) achieves $H_{\infty}=7.859$ bits, only slightly below the $8.092$ bits when using every sensor.
 Parallel findings arise in the PerilZIS and SHL datasets, where omitting a small number of sensors from the “All sensors” set has negligible impact on $H_{\infty}$.
-This pattern appears across datasets: additional modalities may raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
+This pattern appears across datasets: additional modalities raise $H_{0}$ and $H_{1}$ but barely move $H_{\infty}$.
 (a) UCI-HAR
@@
 The results of this analysis are presented in Figure 4.
+For completeness, we also give an alternative binning strategy—Scott’s rule ($3.5\sigma\cdot n^{-1/3}$)—as a comparator for our FD decision rule, which we report as a stability check.
 The graphs show a consistent trend across all datasets.
 For a low number of bins, all entropy measures are severely underestimated as expected.
-As the number of bins increases, the entropy rises sharply before tapering off and converging towards a stable plateau that approximates a logarithmic shape.
-(An exception is the min-entropy value for SHL, which converges after approximately 100 bins).
+As the number of bins increases, the entropy rises sharply before converging towards a stable plateau that approximates a logarithmic shape.
+(An exception is the min-entropy of SHL, which converges after $\approx$100 bins).
 Crucially, FD-selected bin counts fall within the stable region, supporting FD as a balanced choice between fidelity and tractability.
@@
 (d) SHL
-Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set. Freedman–Diaconis (FD) values are also shown.
+Figure 4: Bin sensitivity analysis showing how entropy values vary by average bin width for the best-performing sensor combinations for each data set.
+Freedman–Diaconis (FD) and Scott’s rule values are also shown.
 5 Evaluation
-This section interprets our entropy findings for system design: we discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
+We now discuss the key takeaways, relate min-entropy to concrete guessing cost, discuss mitigations, and give recommendations for future research in the area.
 5.1 Discussion
@@
 Single sensors yield low entropy.
-Building on prior work, we find that min-entropy often remains significantly below the Shannon estimates (e.g.
-some single-sensor readings yield only 1–3 bits).
-Our min-entropy estimates often remain significantly lower than their Shannon counterparts (e.g.
-some single-sensor readings yield only 1–3 bits of min-entropy).
-This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors, which is a critical distinction that existing schemes and evaluations do not capture at this scale.
+Building on prior work, we find that sensor data has stubbornly low min-entropy (e.g.
+some single-sensor readings yield only 1–3 bits) and remains significantly below Shannon entropy.
+This exposes a major gap between average- and worst-case unpredictability across 25 widely used sensors—a critical distinction that existing schemes and evaluations do not capture at this scale.
 Combining multiple sensors offers modest gains with respect to worst-case min-entropy.
-While using multi-modal sensors typically raises best- and average-case entropies ($H_{0}$, $H_{1}$), our experiments reveal that gains in worst-case min-entropy ($H_{\infty}$) are far smaller than one might expect.
+While using multi-modal sensors typically raises best- and average-case entropies, our experiments reveal that gains in worst-case min-entropy are far smaller than one might expect.
 In many cases, $H_{\infty}$ plateaus, reflecting the dominance of a handful of high-probability joint outcomes.
-Even the most complex multi-modal combinations in our datasets yield only modest worst-case entropy, well below the levels required by modern security standards.
+Even the most complex multi-modal combinations in our datasets yield only modest worst-case entropy—between 8–24 bits in the best cases—which is well below the levels required by modern security standards.
 This also aligns with prior work showing that Shannon entropy can significantly overstate security arikan2002inequality , cachin1997entropy , turan2018recommendation , bsi2024ais31 .
@@
 We now sketch how low-entropy sensor data translates into concrete attack cost.
-Let $Z$ be the discretised joint sensor vector used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy.
-Ordering guesses by decreasing probability, the success probability after $q$ guesses satisfies
+Let $Z$ be a random variable over the discretised joint sensor measurement space $\mathcal{Z}$ used by a scheme, and let $H_{\infty}(Z)$ denote its min-entropy.
+If an adversary enumerates candidates $z\in\mathcal{Z}$ in order of decreasing probability $\Pr[Z=z]$, then after $q$ guesses the success probability is bounded by
 $\Pr[\text{success within }q]\;\leq\;q\cdot 2^{-H_{\infty}(Z)}.$ (10)
-The expected number of guesses is $\approx 2^{H_{\infty}(Z)-1}$ cachin1997entropy .
+The expected number of guesses is $E[G]\approx 2^{H_{\infty}(Z)-1}$ cachin1997entropy in order to identify the correct sensor vector.
 Most systems accept similar, not exact, matches through learned decision boundaries or fuzzy matching mehrnezhad2015tap , gurulian2018good , shrestha2014drone , markantonakis2024using , alzubaidi2016authentication , mayrhofer2021adversary , abuhamad2020sensor .
-Let us consider an abstract authentication function, $\mathrm{Auth}(z)\to\{0,1\}$, which accepts or rejects an input vector $z\in Z$.
-An adversary’s task becomes easier through a similarity-based scheme: they only need to generate a candidate $z^{\prime}$ that is ‘close enough’ to a high-probability vector $z$.
-This means the effective success probability for each guess is, in fact, higher than the bound in Eq.
+Let $\mathsf{Auth}:\mathcal{Z}\to\{0,1\}$ be an abstract authentication function that accepts ($1$) or rejects (0) an input vector $z\in\mathcal{Z}$.
+Under such a similarity-based scheme, the adversary does not need to reproduce the exact enrolled sensor vector; it is sufficient to generate some $z^{\prime}$ that lies within the acceptance region of a high-probability template.
+This means the effective success probability for each guess is higher than the bound in Eq.
 10 suggests.
@@
 Indeed, prior work has already demonstrated the feasibility of guessing biometric templates using this approach mai2017guessability .
-In the worst cases, such as the single-sensor readings where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring approximately $2^{7}=128$ guesses on average.
-A main defence against this is to employ rate-limit the number of attempts, although it does not solve the root problem.
-If each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$.
-That is, rate limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
+In the worst cases, such as single-sensor modalities where $H_{\infty}<8$ bits, the search space is trivial for an attacker to exhaust, requiring $\approx 2^{7}$ guesses on average.
+We shall now give more concrete indicative guesswork efforts in line with Cachin cachin1997entropy .
+Consider an adversary who has unlimited guess attempts against $\mathrm{Auth}$, which returns an immediate success/failure result.
+We assume the attacker submits $q$ queries, i.e.
+measurement(s), in descending order of their probability.
+The attacker effort can thus be characterised by the results given in Table 9.
+A primary defence here is to rate-limit the number of attempts; if each guess requires a live interaction, and the system enforces a limit of $r$ attempts per second, the expected time to success is scaled to $\approx 2^{H_{\infty}-1}/r$.
+Nevertheless, such limits linearly stretch attack time but do not change the fundamental feasibility of guessing when $H_{\infty}(Z)$ is in the low-to-mid tens of bits.
+Table 9: Illustrating attacker effort versus min-entropy under optimal online guessing.
+$H_{\infty}$ $E[G]$ (guesses) Time q=1q{=}1/s q=10q{=}10/s $q{=}10^{3}$/s $q{=}10^{6}$/s
+8 $128$ $128$ s $12.8$ s $0.128$ s $0.128$ ms
+12 2,0482{,}048 $34.1$ min $3.41$ min $2.05$ s $2.05$ ms
+16 32,76832{,}768 $9.10$ h $54.6$ min $32.8$ s $3.28$ ms
+20 524,288524{,}288 $6.07$ d $14.6$ h $8.74$ min $0.524$ s
+24 $8.39\times 10^{6}$ $97.1$ d $9.71$ d $2.33$ h $8.39$ s
 5.3 What can be done?
@@
 Quantify effects of deliberate interaction and induced phenomena
-User actions such as prescribed shakes or gestures, as well as externally induced signals, can inject variability and modestly raise entropy by several bits voris2011accelerometers , lv2020analysis .
+User actions such as prescribed shakes or gestures, or externally induced signals, can raise entropy by several bits voris2011accelerometers , lv2020analysis .
 For example, altering the device’s environment to produce distinctive electromagnetic signatures may be detectable by onboard sensors.
@@
 The derivation of shared secrets must come from a component designed for unpredictability; any sensor contribution should be considered, at most, as a small additive input to an entropy pool (see suciu2011unpredictable , wallace2016toward ).
-Cryptographic extractors (e.g.
-Von Neumann and later constructions von195113 , trevisan2001extractors , barak2003true ) or decorrelation techniques (e.g.
-PCA) may be considered to convert low-quality sensor noise into higher-entropy sources.
-However, while extractors reduce bias and correlation, they cannot increase intrinsic min-entropy; if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more.
-Extractors improve the quality of usable randomness when sufficient $H_{\infty}$ already exists; they do not increase the quantity in a way that raises brute-force cost.
-Future work in the area should focus on how entropy can be meaningfully increased while retaining application utility.
+To improve the statistical quality of sensor-derived noise, one may apply cryptographic extractors (e.g.
+the Von Neumann extractor and newer schemes von195113 , trevisan2001extractors , barak2003true ) or decorrelation methods that attenuate dominant, deterministic trends.
+For example, principal component analysis (PCA) has been used elsewhere to isolate high-variance, predictable components—for example, gravitational bias in accelerometer axes—from low-variance residuals dominated by stochastic noise, which may then be discretised into candidate bitstrings (e.g.
+liu2025post ).
+The data processing inequality is relevant here, where $H_{\infty}(f(X))\leq H_{\infty}(X)$, for any deterministic transformation, $f$ vadhan2012pseudorandomness .
+That is, while extractors and transformations can reduce bias and correlation, they cannot increase intrinsic min-entropy; if sensors offer $k$ bits of $H_{\infty}$, post-processing can yield (at best) an unbiased string of length $\approx k$, and no more.
+Thus, such transformations are best interpreted as preconditioning or whitening steps that facilitate later modelling or extraction, but cannot increase the underlying min-entropy of the source itself.
+Future work should focus on designing sensor-based schemes that maximise shared entropy between devices while preserving the practical utility of sensor data for their intended applications.
 5.4 Scope and Limitations
@@
 [^5] By way of example, we could consider a voltage-based fault injection attack that causes nondeterministic errors in the output of individual sensors, affecting the entropy of their output values.
-We also estimate unconditional entropies such that all measurements are independent and identically distributed (i.i.d.).
-This is an optimistic assumption, as sensor data is inherently time-correlated; a sensor’s next reading is often highly dependent on its previous readings.
-For Shannon entropy and for the conditional Rényi notions standard in cryptography (e.g.
-min-entropy), the conditional entropy is always less than or equal to the unconditional entropy; that is, $H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see iwamoto2013information for additional properties).
-Thus, the unconditional entropy values we calculate in this work act as an upper bound.
+Throughout this work,, we report unconditioned, per-sample entropies for each sensor.
+This treats consecutive samples as independent and identically distributed (i.i.d.), which yields an upper bound on unpredictability because real sensor streams are time-correlated.
+For any additional side information, $Y$ (e.g.
+past samples), conditioning cannot increase entropy, i.e.
+$H_{\alpha}(X\,|\,Y)\leq H_{\alpha}(X)$ (see iwamoto2013information for additional properties).
 We may wish to condition $X$ on a sequence of previously observed measurements; however, efficiently quantifying $H_{\alpha}(X_{t}\,|\,X_{<t})$ for high-dimensional sensor streams remains non-trivial.
-Nevertheless, we posit that this would only serve to decrease entropy, strengthening our core conclusion that mobile sensors are insecure data sources for security applications.
+Similarly, we may consider the entropy of a block of measurements over time.
+Under naïve i.i.d.
+conditions, and for intuition purposes only, the entropy rate of a source using block size $N$ and sampling frequency, $f_{s}$, can be considered at most $f_{s}\,\cdot\,N$; for the average modality in PerilZIS (3.912), this naïvely yields a maximum of $\approx$39 bits/sec under a 10Hz sampling frequency.
+Any temporal dependence can only reduce this rate (cf.
+§6, NIST SP 800-90B turan2018recommendation ).
+A full temporal analysis warrants a dedicated study, but we posit that its effects would be to decrease the figures reported in this work, reinforcing our conclusion that mobile sensors are weak sources of unpredictability for security applications.
 6 Conclusion
@@
 Acknowledgments
-Carlton Shepherd has received funding from the UK EPSRC ‘Chameleon’ project (EP/Y030168/1). The authors would like to thank Darren Hurley-Smith for insightful conversations on the topic.
+Carlton Shepherd received funding from the UK’s EPSRC (EP/Y030168/1) and would like to thank Darren Hurley-Smith for insightful conversations on the topic.
+Appendix A Mutual Information Heatmaps
+Figure A.1 shows the pairwise mutual information (in bits) between the sensors of each dataset.
+(a) UCI-HAR
+(b) PerilZIS
+(c) Relay
+(d) SHL
+Figure A.1: Pairwise mutual information (in bits) heatmaps for each dataset.